### v2

# Seeding with Differentially Private Network Information

[…]
Seeding with Differentially Private Network Information
[…]
Seeding with Differentially Private Network Information
[…]
When designing interventions in public health, development, and education, decision makers rely on social network data to target a small number of people, capitalizing on peer effects and social contagion to bring about the most welfare benefits to the population. Developing new methods that are privacy-preserving for network data collection and targeted interventions is critical for designing sustainable public health and development interventions on social networks. In a similar vein, social media platforms rely on network data and information from past diffusions to organize their ad campaign and improve the efficacy of targeted advertising. Ensuring that these network operations do not violate users’ privacy is critical to the sustainability of social media platforms and their ad economies. We study privacy guarantees for influence maximization algorithms when the social network is unknown, and the inputs are samples of prior influence cascades that are collected at random. Building on recent results that address seeding with costly network information, our privacy-preserving algorithms introduce randomization in the collected data or the algorithm output, and can bound each node’s (or group of nodes’) privacy loss in deciding whether or not their data should be included in the algorithm input. We provide theoretical guarantees of the seeding performance with a limited sample size subject to differential privacy budgets in both central and local privacy regimes. Simulations on synthetic and empirical network datasets reveal the diminishing value of network information with decreasing privacy budget in both regimes.
[…]
social network interventions, differential privacy, influence maximization, sample complexity, approximation algorithms, peer influence and social contagion, public health interventions
1 Introduction
When designing interventions in public health, development, and education decision-makers rely on social network data to target a small number of people, capitalizing on peer effects and social contagion to bring about the most welfare benefits to the population. Many of these seeding methods are implemented without complete knowledge of social network structures — e.g., by targeting nominated friends of random individuals (Kim et al. 2015) or random pairs of friends (Alexander et al. 2022) — and they enhance adoption of public health interventions by driving indirect treatment effects among people who were affected by spillovers from others’ treatment (even though they were not part of the initial treatment group). However, even without complete network information, network participants (and in particular minority members) who contribute to partial social network data collection endure privacy risks. This is especially important when engaging with vulnerable populations to record sensitive or stigmatizing behavior, e.g., in design of public health interventions among the homeless youth for HIV or suicide prevention (Wilder et al. 2021). Developing new methods that are privacy-preserving for network data collection and targeted interventions is critical for responsible design of public health interventions on social networks and maintaining community engagements that sustain such interventions.
In this paper we propose randomization techniques based on the differential privacy formalism to provide social network participants with plausible deniability. Under these schemes the network participants can make a credible claim that their associates are targeted (e.g., recruited into treatment) by random choice, irrespective of their associations. We investigate how the targeting efficacy is affected by decreasing the privacy budget that entails more noise in the collected data (with local privacy mechanisms) or the targeting algorithm’s output (with central privacy mechanisms).
After summarizing our main contributions in Section 1.1 and providing detailed comparisons with related works in Section 1.2, in Section 2 we provide preliminaries including hardness results that demonstrate the inadequacy of mainstream approaches to social network data privacy for seeding interventions and motivate our proposed notion of “influence sample differential privacy”. In Section 3, we provide privacy mechanisms for seeding in both central (Section 3.1) and local (Section 3.2) regimes. In Section 4, we discuss and contextualize the nature of our privacy protections for social network data collection and intervention design applications. In Section 5, we test the performance of our privacy mechanisms on synthetic and real social network datasets. We end with a discussion of limitations, future directions and concluding remarks in Section 6. Complete proofs and supplementary results are provided in several appendices at the end.
1.1 Summary of the main results
Differential privacy (DP) offers an information-theoretic foundation for privacy by focusing on protecting input data against revelations of algorithmic outputs (Dwork et al. 2014). In mathematical formulation, a randomized mechanism $\mathcal{M}$ is $\epsilon$-DP if for any two adjacent input datasets $D$ and $D^{\prime}$ in its input domain and any subset $A$ in its output range we have:
$\frac{\Pr[\mathcal{M}(D)\in A]}{\Pr[\mathcal{M}(D^{\prime})\in A]}\leq e^{\epsilon}.$ (1)
The $\epsilon$-DP criterion in Equation 1 limits inferences of any observer/adversary about the input data based on the algorithm output. The adjacency notion between datasets is typically defined such that neighboring datasets differ in contributions of an individual entity, and thus $\epsilon$-DP protects individuals’ contributions from ramification of an adversarial inference. Hence, different notions of adjacency will imply different protections for the individuals. In the sequel, we explain our proposed notion of adjacency for seeding with randomized contagion data (which we term “influence samples”), and the nature of privacy protections that it offers. Throughout this paper we use DP to signify any of the phrases “differentially private”, “differentially privately”, and “differential privacy” with or without hyphen between the words as appropriate by the language function that these words play in the context that DP is used.
Seeding with partial and private network information.
We study a classical, NP-hard network optimization problem, known as “influence maximization”, which asks to select $k$ nodes on a graph of size $n$ such that the expected spread size from seeding those $k$ nodes is maximum under a randomized model of network diffusion, whereby each edge passes on the influence (or “infection”) with an independent cascade probability $p$. Given the social network graph ${\mathcal{G}}$ and an integer $k\geq 1$, we denote the expected spread size from infecting the nodes in $S$ by $I_{\mathcal{G}}(S)$ and write $\mbox{\sf OPT}=\max_{S:|S|\leq k}I_{\mathcal{G}}(S)$ for the optimum solution to the $k$-influence maximization problem. It is an example of a cardinality-constrained combinatorial optimization on a graph with a monotone, submodular objective that admits a tight $(1-1/e)$ approximation guarantee using the greedy node selection algorithm. This problem has occupied a central place in graph and social networks algorithms literature since its introduction by Kempe et al. (2003). There is a wealth of literature (spanning almost two decades) on efficient approximations of this problem (Kempe et al. 2003, 2005, 2015, Chen et al. 2009, Wang et al. 2012, Borgs et al. 2014). However, the graph structure of the input data has stunned the design of DP influence maximization algorithms. Common operationalizations of DP for graph algorithms, such as node- or edge-DP define adjacent inputs in Equation 1 based on neighboring graphs differing in a node or an edge. This adjacency criteria is often too stringent for graph problems whose solutions are extremely sensitive to existence of nodes and edges (see Section 1.2 for detailed citations and comparisons). Our main hardness results indicate that it is impossible to obtain useful performance guarantees for influence maximization subject to node- or edge-DP constraints (see Section 2.2, LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg, summarized in 1 below).
Informal Main Results 1
Given $n\geq 1$ and $\epsilon>0$, there can be no $\epsilon$-node-DP or $\epsilon$-edge-DP seeding algorithms that give an approximation guarantee better than $(1-1/e)\mbox{\sf OPT}-\alpha_{\epsilon}{n}$ for $\alpha_{\epsilon}\geq\frac{1}{1000e^{\epsilon}}$.
A new privacy notion for social network data collection and intervention design.
Our main contribution here is to formulate a new differential privacy notion with guarantees for influence maximization when the social network is unknown and the algorithm inputs are samples of prior influence cascades that are collected at random, i.e., “influence samples” — rather than the entire social network graph. Our results build on recent work about seeding with costly network information (Eckles et al. 2022), and more broadly sample complexity of influence maximization (Sadeh et al. 2020, Borgs et al. 2014). Our algorithm input is a set of $m$ influence samples, i.e., $m$ independent cascade samples that are collected on the social network graph each starting from a random, initially-infected node. We denote this input deadset by a binary matrix ${\mathbf{x}}\in\{0,1\}^{n\times m}$, where each row represents a node and each column represents the status of the node in an influence sample: $[{\mathbf{x}}]_{ij}=1$ if node $i$ is infected in the $j$-th influence sample and $[{\mathbf{x}}]_{ij}=0$ otherwise. Our privacy notion is with respect to the contribution of individual network participants to the influence samples, hence, we define the adjacent input datasets in Equation 1 as ${\mathbf{x}},{\mathbf{x}}^{\prime}\in\{0,1\}^{n\times m}$ which differ in a single binary entry:
Influence Sample Differential Privacy (ISDP)We define “$\epsilon$ influence sample differential privacy ($\epsilon$-ISDP)” by controlling the odds ratio of the output of the seeding algorithm on adjacent influence samples, i.e., those differing in participation of a single node in a single influence cascade.
We follow the formulation of influence maximization by Eckles et al. (2022) where the social network is unknown and influential nodes need to be estimated from costly samples of past spreads. Our privacy guarantees are with respect to influence samples, controlling information leakage about the cascade data. For example, Apple Inc. uses a randomized response mechanism to collect the users’ daily usage of emojis with a fixed privacy budget ($\epsilon=4$). This setup maps naturally to our formulation of ISDP as the usage of each emoji can be modeled as an influence sample. We make no assumptions about adversary’s knowledge of the network and focus on differential privacy guarantees for influence maximization according to the ISDP formulation above.
We analyze both central and local privacy under the ISDP framework. In central differential privacy (CDP), the mechanism designer is trustworthy and has direct access to the private data (influence samples). For local differential privacy (LDP), the data can be aggregated by some (not necessarily trustworthy) third party. Our CDP mechanism only protects (adds noise to) the output of seed selection. In contrast, our LDP mechanism adds noise to the influence samples directly.
Sample complexity with differential privacy and approximation guarantees.
By introducing randomization at the input or output of seeding algorithms, we are able to bound each node’s privacy loss in deciding whether or not their data should be included in the influence samples — as opposed to the social network graph. Of note, this is critically different from DP analysis of graphs that is defined with respect to neighboring graphs in the algorithm input. In our approach, each node’s decision to appear in the input does not affect the data generation process (DGP) of the input sample or the underlying diffusion process; the network graph remains the same and the privacy implications are studied with respect to the construction of the samples on a fixed network. This allows us to meaningfully bound the effect of the removal of a node’s data on the algorithmic performance and to use those bounds to tradeoff the privacy risks against improved performance. To mitigate privacy risks, DP uses randomization to provide plausible deniability certificates to those contributing to the input data. This is achieved in one of the two ways: (i) using exponential mechanism (McSherry and Talwar 2007) by randomizing the algorithm output, i.e., central DP, (ii) using the randomized response mechanism (Warner 1965) by injecting noise in the input data, i.e., local DP. Accordingly, we propose two efficient DP seeding algorithms and bound their performance in terms of the sample size $m$, network size $n$, seed set size $k$, and privacy budget $\epsilon$, (See Section 3 LABEL:thm:exp:acc and LABEL:thm:rr:acc):
Informal Main Results 2
Consider a graph ${\mathcal{G}}$ on $n\geq 2$ nodes and $k\leq n$, for any $0<\alpha\leq 1$ and $\epsilon>0$, we give a set of algorithms that are centrally or locally $\epsilon$-ISDP and their outputs infect at least $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes with high probability. Our central $\epsilon$-ISDP algorithm uses exponential mechanism and requires at least $m=\max\{\frac{12}{\alpha\epsilon},\frac{9}{\alpha^{2}}\}k\ln n$ influence samples and runs in $O(knm)$ time. Our local $\epsilon$-ISDP algorithm uses randomized response to flip each node’s status in the influence samples with independent probabilities $\rho_{\epsilon}=1/(1+e^{\epsilon})$. Compared to our central DP seeding algorithm, the local algorithm requires more (exponentially in $k$) influence samples: $m=O({k^{3}V_{\rho_{\epsilon},k}^{2}\ln n}/{\alpha^{2}})$ with $O(nk^{4}+kn^{2}m)$ run time, where $V_{\rho_{\epsilon},k}=O(\max\{2,1/\epsilon\}^{k^{2}})$.
Given $m$ influence samples, in Section 3.1 we use exponential mechanisms (McSherry and Talwar 2007) to choose high quality seeds with central DP guarantees. In Subsection 3.2, before accessing the data (influence samples), we use the randomized response mechanism (Warner 1965), to perturb the input data and protect privacy at the local level. However, these perturbations can introduce additional biased noise into the data, so we provide a closed-form characterization of the introduced noise and an algorithm to offset this bias. In Section 5, we thoroughly evaluate our algorithms on synthetic and real graphs of different sizes and verify our theoretical guarantees in simulations.
1.2 Comparisons with related work
Seeding network diffusions.
A long line of literature, going back to the seminal works of Kempe et al. (2003, 2005, 2015) propose efficient approximation algorithms for this problem based on the submodularity of the influence function (Chen et al. 2009, Wang et al. 2012). This problem arise naturally in applications such as development economics and public health to facilitate adoption of welfare-enhancing technologies, e.g., microfinance (Banerjee et al. 2013), insurance (Cai et al. 2015), contraception (Behrman et al. 2002), water purification methods, and multivitamins (Kim et al. 2015); and to target localized social network interventions for spreading information about immunization camps (Banerjee et al. 2019), preventing misinformation about drug side effects (Chami et al. 2017), increasing HIV awareness among homeless youth (Yadav et al. 2017), and reducing bullying among adolescent youth (Paluck et al. 2016). Alexander et al. (2022) show that network-based targeting substantially increases population-level adoption of iron-fortified salt to prevent iron-deficiency anemia among female household heads in India. Kim et al. (2015) show similar results for a cluster-randomized trial in villages in Honduras, where they measure the efficacy of social network targeting strategies to promote adoption of chlorine for water purification and multivitamins for micronutrient deficiencies. In all such settings, collecting complete social network connection (edge) data is difficult, costly, or impossible and comes with critical privacy and security risks to individuals in the social network. In this paper we focus on differential privacy guarantees for influence maximization when social network is unknown and influential nodes need to be estimated from costly samples of past spreads.
Social network data privacy.
Much of the past work on social network data privacy focuses on identifiability of nodes from released graph data, with or without additional information on node attributes and their neighborhoods (Romanini et al. 2021). Classical anonymization techniques mask node attributes and perturb, modify, randomize or aggregate the graph structure to prevent re-identification of nodes within a confidence level — typically achieving k-anonymity (Wu et al. 2010, Zheleva and Getoor 2011, Abawajy et al. 2016). However, statistical disclosure control of collected data requires safeguarding data donor’s privacy against adversarial attacks where common anonymization techniques are shown to be vulnerable to various kinds of identification (Barbaro et al. 2006), linkage and cross-referencing (Narayanan and Shmatikov 2008, Sweeney 2015, 1997), statistical difference and re-identification attacks (Kumar et al. 2007). On the other hand while edge sampling at the data collection can somewhat mitigate the re-identification risk of nodes (Romanini et al. 2021, Jiang et al. 2021), the seeded nodes can reveal additional side information about who has contributed to network data collection. For example, once a community leader is chosen, his or her close contacts may suffer privacy leaks, even though no social network data are published. Our privacy guarantees for influence maximization follow the DP formalism to provide network participants and data donors with plausible deniability with regard to their contribution to the collected data or their role in shaping the algorithm output (the seeded nodes). DP is a gold standard for privacy-preserving data analysis and avoids the above drawbacks by abstracting away from the attack models and controlling the privacy leak at an information-theoretic level according to Equation 1.
DP analysis of graphs.
Differential privacy (DP) is a versatile framework for developing privacy-preserving algorithms, and it has been impactful in large-scale and high-profile applications such as the U.S. decennial census Garfinkel (2022) and the tech industry Erlingsson (2014a), Apple Differential Privacy Team (2017). It have been successfully applied to a broad variety of problem classes, including, e.g., combinatorial (Gupta et al. 2010), submodular (Mitrovic et al. 2017, Sadeghi and Fazel 2021), and online submodular (Salazar and Cummings 2021) optimization problems. The existing literature on DP analyses of graphs, e.g., edge-DP Nissim et al. (2007) or node-DP (Kasiviswanathan et al. 2013, Raskhodnikova and Smith 2016), has devoted much attention to graph statistics and parameters such as average degree and various subgraph counts (Blocki et al. 2022), degree distribution (Hay et al. 2009) and the graph spectra (Wang et al. 2013), as well as community Mohamed et al. (2022) and densest subgraph detection (Nguyen and Vullikanti 2021). These techniques are not directly applicable to influence maximization, and our main hardness results indicate that node- or edge-DP are too stringent to render useful performance guarantees for influence maximization. Similar observations have been made about node- and edge-DP guarantees of other graph algorithms which are too sensitive to the presence of individual edges (and especially nodes) in their input graphs (Gupta et al. 2010, Li et al. 2022, Epasto et al. 2022, Blocki et al. 2013). In their analysis of deferentially private combinatorial optimization, Gupta et al. (2010) consider several graph problems including vertex and set cover. Their modified greedy algorithm for set cover is similar to our central DP mechanism in Section 3.1, though addressing a substantially different privacy environment, where the goal is to protect a private subset $R$ while outputting a minimal collection that covers $R$. In both cases direct application of the exponential mechanism is inefficient and the greedy approach facilitates an efficient implementation.
In formulating our notion of “influence sample differential privacy” (ISDP), we move away from having the social network as the algorithm input. Instead, we focus on an input dataset of $m$ cascades that are collected at random on the social network graph. ISDP allows us to limit the information leakage about contribution of individual nodes to the input cascades (with local ISDP), or contribution of individual nodes to the algorithm output through the input cascades (with central ISDP). It is worth noting that achieving local privacy under ISDP is easy by randomizing individual statuses of the nodes in the input cascades. However, in general it is challenging to coordinate randomization across nodes in a meaningfully decentralized manner that matches the problem context, and locally deferentially private analysis of graphs has been limited so far to certain restricted settings (Imola et al. 2021, Dhulipala et al. 2022).
Privacy of network diffusions.
The fundamental relationship between information flow and privacy is well noted, cf., contextual integrity (Nissenbaum 2009, Benthall and Cummings 2022). Notwithstanding, the privacy implications of information diffusion and algorithmic intervention on social networks is largely unexplored, with the exception of two studies (Rezaei et al. 2021, Liell-Cock et al. 2020). The existing work looks at this issue in specific, highly-stylized scenarios. One study shows that it is extremely hard to hide the existence of a giant, connected component of infected nodes under an independent cascade model and a simple inference attack can reveal the status of a good fraction of nodes (Rezaei et al. 2021). In another study, the authors propose a decaying pseudo-noise that is added locally to mask the influence structure against an external observer with access to the entire information flow under the Friedkin-Johnsen influence propagation model (Liell-Cock et al. 2020). With our formulation of DP for influence maximization we can focus on privacy leaks through information flow on the network rather than the underlying network structure. Our formulation of DP for influence samples ($\epsilon$-ISDP in 1) allows us to explicitly quantify information leaks from observation of random spreads on the network and for us to take a direct approach to the privacy risks that social network participants face from network diffusions.
2 Problem Setup and Preliminary Results
2.1 Influence sampling and maximization
Independent cascade model.
In the independent cascade model, infections spread via an edge-weighted undirected graph ${\mathcal{G}}=(V,p)$ on $n$ nodes. The set of nodes, $V$, is indexed by $[n]:=\{1,\ldots,n\}$ and $p:V\times V\to[0,1]$ assigns probability weights to every edge in ${\mathcal{G}}$. An infection begins at a set $S$ of seed nodes and spreads through the graph in rounds. Each infected node, once they become infected, has a single chance of passing the infection to their neighbors. Each undirected edge $e=(v,u)$ has a weight $p_{e}\in[0,1]$ representing the probability that the process spreads along edge $e$ to node $u$ in the round immediately after $v$ is first infected.
The above process has the following equivalent formulation: We first sample an undirected, unweighted graph $g$ by independently realizing each edge $e$ with probability $p_{e}$. We write $g\sim{\mathcal{G}}$ to denote that $g$ is drawn from the random graph distribution ${\mathcal{G}}$. Then we can construct the set of infected nodes in the independent cascade process, as the set of nodes that are reachable from seed set $S$ in $g$. Now we define some notations. Given an undirected graph $g$ and subsets of nodes $S$ and $T$, we write $C_{g}(S)$ for the set of nodes that are reachable from (or infected by) $S$ in $g$, and $C_{g}(S|T):=C_{g}(S)\setminus C_{g}(T)$ for the set of nodes that are reachable from $S$ but not from $T$. Then the influence function $I_{g}(S):=|C_{g}(S)|$ is the number of nodes reachable (or infected) by $S$ and $I_{\mathcal{G}}(S):=\operatorname*{\mathbb{E}}_{g\sim{\mathcal{G}}}[I_{g}(S)]$ is its expectation. Similarly, we define the conditional influence function $I_{g}(S|T):=|C_{g}(S|T)|$ and denote its expected value by $I_{\mathcal{G}}(S|T):=\operatorname*{\mathbb{E}}_{g\sim{\mathcal{G}}}[I_{g}(S|T)]$.
The influence maximization problem.
Given graph ${\mathcal{G}}$ and integer $k\geq 1$, the $k$-influence maximization problem is to find a set $S\subset V$ of at most $k$ nodes maximizing the expected number of reachable nodes, $I_{\mathcal{G}}(S)$. We write $\mbox{\sf OPT}=\max_{S:|S|\leq k}I_{\mathcal{G}}(S)$ for the maximum expected influence of any set of $k$ nodes.
[…]
Influence sample greedy seeding algorithm.
How can we design a seeding algorithm $M$ that takes $m$ influence samples ${\mathbf{x}}$ and outputs a set of $k$ nodes, $M({\mathbf{x}})\in\mathcal{Y}:=\{0,1\}^{V}$? Here we present a seeding algorithm by Eckles et al. (2022). The core idea of Algorithm 1 is to estimate the influence function from the influence samples, and run the greedy algorithm. Specifically, given $m$ realized influence samples ${\mathbf{x}}$ and a subset of nodes $S$ in ${\mathcal{G}}$, we define the $S$-reachable influence samples as: $C_{\mathbf{x}}(S):=\{t\in[m]:\exists v\in S,x_{v}^{t}=1\},$ which is the set of influence samples whose activated nodes are reachable from the set $S$, and $C_{\mathbf{x}}(S|T)=C_{\mathbf{x}}(S)\setminus C_{\mathbf{x}}(T)$. Alternatively, if we treat ${\mathbf{x}}$ a collection of $m$ subset $x^{1},\dots,x^{m}$, $C_{\mathbf{x}}(S)=\{t:S\cap x^{t}\neq\emptyset\}$. Finally, we define $I_{\mathbf{x}}(S)=\frac{n}{m}|C_{\mathbf{x}}(S)|$ and $I_{\mathbf{x}}(S|T)=\frac{n}{m}|C_{\mathbf{x}}(S|T)|$. The following proposition says the expectation of $I_{\mathbf{x}}$ is equal to $I_{{\mathcal{G}}}$. We will use $v$ as $\{v\}$ for singleton sets to simplify the notation, e.g., $I_{\mathcal{G}}(v)$ as $I_{\mathcal{G}}(\{v\})$.
Proposition 2.1 (Observation 3.2. of Borgs et al. (2012) – cf. Borgs et al. (2014))
Given a positive integer $m$, a graph ${\mathcal{G}}=(V,p)$ and sets of nodes $S,T\subseteq V$, if ${\mathbf{x}}$ is the random variable of $m$ influence sample, then $\operatorname*{\mathbb{E}}_{{\mathbf{x}}}[I_{{\mathbf{x}}}(S)]=I_{\mathcal{G}}(S)$, and $\operatorname*{\mathbb{E}}_{\mathbf{x}}[I_{\mathbf{x}}(S|T)]=I_{\mathcal{G}}(S|T)$. Additionally, by Hoeffding’s inequality, for all $\epsilon>0$ we have: $\Pr[|I_{{\mathbf{x}}}(S)-I_{\mathcal{G}}(S)|\geq\epsilon n]\leq 2\exp(-2\epsilon^{2}m),$ where the probability is over the randomness in the generative model of influence samples $x$.
With the above observation, Eckles et al. (2022) design a seeding algorithm that only uses influence samples and show that if the number of influence sample is of order $k^{2}(\log n/\alpha)/\alpha^{3}$, then Algorithm 1 is optimal up to $\alpha n$ additive error. Informally, Algorithm 1 is a greedy algorithm that sequentially query Algorithm 2 for the seed that maximizes the marginal improvement. In this paper, we will follow similar idea that combine greedy algorithm and an estimation algorithm for marginal improvement.
Algorithm 1 Influence sample greedy seeding algorithm
1:$m$ influence samples ${\mathbf{x}}$ and total number of seeds $k$.
2:Set $S_{0}=\emptyset$
3:for $i$ from $1$ to $k$ do
4:    $v_{i}=\mathcal{M}({\mathbf{x}},S_{i-1})$ by Algorithm 2
5:    $S_{i}=S_{i-1}\cup v_{i}$
6:end for
7:Return $S_{k}$
Algorithm 2 Single seed selection, $\mathcal{M}({\mathbf{x}},S)$
1:$m$ influence samples ${\mathbf{x}}$ and a seed set $S$.
2:for $v\in[n]\setminus S$ do $\triangleright$ Chooses the $|S|+1$-th seed.
3:    Compute $I_{\mathbf{x}}(v|S)=\frac{n}{m}|C_{\mathbf{x}}(S\cup v)\setminus C_{\mathbf{x}}(S)|$.
4:end for
[…]
In Algorithm 1, we slightly modify the INF-SAMPLE$(m,k)$ algorithm to sequentially call Algorithm 2 for selecting the next seed. Instead of collecting fresh influence samples at each greedy step, Algorithm 1 calls Algorithm 2 to reuse the influence samples and have much better influence sample complexity as a result. We formally show this improvement in LABEL:thm:exp:acc.
Theorem 2.2 (label = thm:original, name = Theorem 2 of Eckles et al. (2022))
Given ${\mathcal{G}}$ of $n\geq 2$ nodes and $k\leq n$, for any $0<\alpha\leq 1$, the output of Algorithm 1 can infect $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes using no more than $81k^{2}\log(6nk/\alpha)/\alpha^{3}$ influence samples.
It is well known that even with the full knowledge of the graph ${\mathcal{G}}$ no efficient algorithm can perform better than $(1-1/e)\mbox{\sf OPT}$ unless P equals NP. Eckles et al. (2022) show that without the knowledge of ${\mathcal{G}}$, Algorithm 1 can approach an optimal efficient algorithm when the number of influence samples is large enough. In this paper, we ask if we can achieve similar performance when privacy constraints are imposed. Eckles et al. (2022) also show that one cannot hope to provide a constant factor approximation guarantee using $o(n)$ influence samples.
Theorem 2.3 (Theorem 3 of Eckles et al. (2022))
Let $10<\alpha<$ 1 be any constant. For any algorithm with $o(n)$ samples, there exists a weighted graph ${\mathcal{G}}$ so that the output of the algorithm can only infect less than $\alpha\mbox{\sf OPT}$ nodes.
2.2 Influence Sample Differential Privacy
Intuitively, a differentially private (DP) algorithm ensures that the algorithm outputs on two adjacent inputs are close to each other, making them hard to distinguish. However, there are multiple alternative definitions of adjacency relation on graph-structured input data. One classical notion is node DP where two graphs are adjacent if they are different by one node’s connections. Another one is edge DP where two graphs are adjacent if they differ by one edge (Hay et al. 2009). In this paper, we would like to focus on DP seeding algorithms that use influence samples but not raw graph data. Instead of raw graph, we study algorithms that only use historical cascades data — i.e., influence samples — where two collections of $m$ influences samples ${\mathbf{x}}$ and ${\mathbf{x}}^{\prime}$ in $\{0,1\}^{n\times m}$ are (influence sample) adjacent if we can change one entry to convert $\mathbf{x}$ to $\mathbf{x}^{\prime}$, and write ${\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}$ if they are adjacent. Given the above adjacency relation on influence samples, we can define if a stochastic seeding algorithm $M$ is DP.
Definition 2.4 (ISDP)
Given $\epsilon\geq 0$, $n,m\geq 1$, and $k\in[n]$, a function $M^{k}:\{0,1\}^{n\times m}\to\mathcal{Y}$ is $\epsilon$-influence sample differentially private ($\epsilon$-ISDP, in short) if for all outputs ${\mathbf{y}}\in\mathcal{Y}$ of $k$ seeds, and all pairs of adjacent datasets (collection of influence samples), $\mathbf{x}\underset{is}{\sim}\mathbf{x}^{\prime}$, we have: $\Pr[M^{k}(\mathbf{x})={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}(\mathbf{x}^{\prime})={\mathbf{y}}].$
Instead of entry-level adjacency, one may consider individual-level adjacency that two inputs are adjacent if they differ at one row (i.e., an individual can decide whether to report their whole history or not). Our analysis in the paper still holds after strengthening the privacy constraints from $\epsilon$ to $\epsilon/m$; please refer to Section 6 for a more detailed discussion.
We believe the study of influence sample DP is more natural and practical than conventional node and edge DP for seeding algorithms (Kasiviswanathan et al. 2013, Raskhodnikova and Smith 2016). First, the notion of node DP allows an agent to misreport her connection in the social networks. However, when collecting data on network cascades, a single agent can hardly modify her connection to change the dynamics of social contagion, but rather she can decide whether to misreport her previous contagion data. Therefore, we argue that defining adjacent datasets based on influence samples as the input, rather than the raw social network data, is more natural and realistic. Moreover, we show in LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg that node-DP and edge-DP are too strict for influence maximization and hardly admit any utility guarantees. In Section 4 we will elaborate on the nature of our privacy protections in social network data collection and intervention design contexts.
Give two edge-weighted undirected graphs ${\mathcal{G}}=(V,p)$ and ${\mathcal{G}}^{\prime}=(V^{\prime},p^{\prime})$, we say ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ are node adjacent if $V=V^{\prime}$ and there exists $v\in V$, $p_{e}=p_{e}^{\prime}$ for all $v\notin e$. Similarly, ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ are edge adjacent if $V=V^{\prime}$ and there exists $u,v\in V$ $p_{e}=p_{e}^{\prime}$ for all $e\neq(u,v)$. For $k<n$, we say that a seeding algorithm $M^{k}$ is $\epsilon$-node (or edge) DP if for any $k$, node (or edge) adjacent graphs ${\mathcal{G}},{\mathcal{G}}^{\prime}$, and output ${\mathbf{y}}$, $\Pr[M^{k}({\mathcal{G}})={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}({\mathcal{G}}^{\prime})={\mathbf{y}}]$.
The following proposition (proved in Section 7.1) shows no node DP seeding algorithm can output a high quality seed.
Proposition 2.5 (label = prop:node:dp:neg, name=Hardness of Influence Maximization with Node-DP)
Given $n\geq 1$ and $\epsilon>0$, for any $\epsilon$-node-DP seeding algorithm $M^{k}({\mathcal{G}})$, there exists ${\mathcal{G}}$ of $n$ nodes so that for $k=1$ we have: $\max_{S:|S|=1}I_{\mathcal{G}}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}({\mathcal{G}})}[I_{\mathcal{G}}({\mathbf{y}})]\geq\frac{n}{4}-(e^{\epsilon}+1).$
We use the fact that a star and a totally isolated graph are different by one node’s connection. A node-DP algorithm needs to output similar seed sets on these two graphs, and thus returns low quality seeds. Based on similar intuition, the following proposition (proved in Section 7.2) shows no edge-DP seeding algorithm can output a high quality seed either. Both results can be strengthened to exclude the possibility of useful algorithms with “approximate” $(\epsilon,\delta)$-DP guarantees as well. The latter is a relaxation of the defining constraint in Equation 1 to allow for an additive term $\delta$; see Section 7.2.
Proposition 2.6 (label = prop:edge:dp:neg, name=Hardness of Influence Maximization with Edge-DP)
For all $n\geq 1$ and $\epsilon>0$, there exists $\alpha_{\epsilon}\geq\frac{1}{1000e^{\epsilon}}$ so that any $\epsilon$-edge-DP seeding algorithm $M^{k}({\mathcal{G}})$, there exists graph ${\mathcal{G}}$ on $n$ nodes so that for $k=1$ we have: $(1-1/e)\textstyle\max_{S:|S|=1}I_{\mathcal{G}}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}({\mathcal{G}})}[I_{\mathcal{G}}({\mathbf{y}})]\geq\alpha_{\epsilon}n.$
Compared to LABEL:thm:original, the above results show that no node-DP (or edge-DP) algorithm can perform close to $(1-1/e)\mbox{\sf OPT}$ even with the full knowledge of ${\mathcal{G}}$. In Section 7.2 (Proposition 7.1), we show that similar lower bounds hold even if we consider an additive relaxation of the node- and edge-DP requirements, allowing the $\Pr[M^{k}({\mathcal{G}})={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}({\mathcal{G}}^{\prime})={\mathbf{y}}]$ condition to be violated up to a $\delta$ additive factor. In this paper, we will design influence sample DP algorithms that perform comparable to Algorithm 1. Finally, note that node-DP implies edge-DP, but influence sample DP is not comparable to node- or edge-DP because the algorithms take different inputs.
Two primitive DP algorithms.
Our algorithms will make use of the following two general DP mechanisms (we use algorithm and mechanism interchangeably): exponential mechanism (McSherry and Talwar 2007, Dwork et al. 2014), and randomized respond mechanism (Warner 1965, Evfimievski et al. 2003). First, to achieve $\epsilon$-DP, given a finite set (range), $\mathcal{R}$, the exponential mechanism is defined with respect to a scoring function $u:\{0,1\}^{n\times m}\times\mathcal{R}\to{\mathbb{R}}$ and its sensitivity $\Delta(u):=\max_{r\in\mathcal{R},{\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}}|u({\mathbf{x}},r)-u({\mathbf{x}}^{\prime},r)|$. The exponential mechanism $\mathcal{M}^{\epsilon}_{E}({\mathbf{x}},u,\mathcal{R})$ outputs an element $r\in\mathcal{R}$ with probability proportional to $\exp\left(\frac{\epsilon u({\mathbf{x}},r)}{2\Delta(u)}\right)$ as defined in Dwork et al. (2014). Moreover, as shown in the following theorem, for a fixed dataset ${\mathbf{x}}$ exponential mechanism’s output almost maximizes the scoring function $u({\mathbf{x}},r)$ and can be interpreted as a DP soft max.
Theorem 2.7 (label = thm:exp:utility, name = Theorem 3.11 of Dwork et al. (2014) — Exponential Mechanism)
Consider a dataset ${\mathbf{x}}\in\{0,1\}^{n\times m}$, privacy budget $\epsilon>0$, a finite set $\mathcal{R}$, and a scoring function $u:\{0,1\}^{n\times m}\times\mathcal{R}\to{\mathbb{R}}$ with sensitivity $\Delta(u)$. The exponential mechanism $\mathcal{M}^{\epsilon}_{E}({\mathbf{x}},u,\mathcal{R})$ that outputs an element $r\in\mathcal{R}$ with probability proportional to $\exp\left(\frac{\epsilon u({\mathbf{x}},r)}{2\Delta(u)}\right)$ is $\epsilon$-ISDP and for all $t>0$, we have:
$\Pr\left[u({\mathbf{x}},\mathcal{M}^{\epsilon}_{E}({\mathbf{x}},u,\mathcal{R}))\leq\max_{r\in\mathcal{R}}u({\mathbf{x}},r)-\frac{2\Delta(u)}{\epsilon}(\ln|\mathcal{R}|+t)\right]\leq e^{-t},$
where the probability is over the randomness of the algorithm run for fixed ${\mathbf{x}}$.
The above exponential mechanism ensures that the output does not reveal too much information about the privacy-critical input data, limiting how much an adversary can infer about the input data according to Definition 2.4. However, depending on data curation requirements (who can access the exact user data), protection against output-based inferences may not be sufficient and we may need to limit the algorithm’s access to exact input data. Such privacy guarantees can be formulated as local DP where each individual’s data is privatized before running the algorithms. One standard locally DP algorithm is the randomized response mechanism (Evfimievski et al. 2003). Specifically, an $\epsilon$-DP randomized response mechanism guarantees privacy by independently flipping each bit of the input dataset (influence sample collection) with probability $1/(e^{\epsilon}+1)$.
3 Differentially Private Seeding Algorithms
In this section, we provide two DP seeding algorithms that satisfy the DP requirements either centrally or locally. First in Subsection 3.1, we use an exponential mechanism that ensures the output of the seed selection is centrally DP. Next in Subsection 3.2, we use the randomized response algorithm to achieve local DP according to the criteria set forth in Definition 2.4. LABEL:thm:exp:acc and LABEL:thm:rr:acc provide our sample complexity bounds in each case to achieve $\epsilon$-ISDP with a $(1-1/e)\mbox{\sf OPT}-\alpha n$ performance guarantee. In particular, we observe that the sample complexity bound for achieving local DP is much harsher at $O(k^{3}\epsilon^{-2k^{2}}\ln n/\alpha^{2})$ compared to $O(k\ln n/(\alpha\epsilon))$ for central DP. This exponential separation of sample complexity dependence on $k$ is reflected in our simulation results in Section 5, where exponential mechanism shows a faster improvement in performance with increasing number of influence samples and its output has a significantly higher expected spread size when the privacy budget ($\epsilon$) is low.
3.1 Central DP with Exponential Mechanism
In this section, we design a DP seeding algorithm by using exponential mechanism and modifying Algorithm 1. To protect privacy, instead of Algorithm 2, we use an exponential mechanism that differentially outputs a soft max (LABEL:thm:exp:utility). Formally, given $\epsilon>0$ and integer $k$ we define a seeding algorithm $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ that runs Algorithm 1 but replaces the single seed selection steps of Algorithm 2 by its exponential mechanism variant, Algorithm 3 $\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}},S)$. Note that Algorithm 3 is an exponential mechanism using $I_{\mathbf{x}}(\cdot|S)$ as the scoring function.
Algorithm 3 $\epsilon$-ISDP single seed selection, $\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}},S)$
1:$m$ influence samples ${\mathbf{x}}=(x^{1},\dots,x^{m})^{\top}$, a seed set $S$, and privacy budget $\epsilon\geq 0$.
2:for $v\in[n]\setminus S$ do
3:    Compute $I_{\mathbf{x}}(v|S)$
4:end for
5:Return $v$ with probability proportional to $\exp\left(\frac{\epsilon m}{2n}I_{\mathbf{x}}(v|S)\right)$
Note that a naive algorithm that runs the exponential mechanism directly on all $k$ seeds at once is not efficient because the size of the output space is combinatorially large $\Omega(n^{k})$. To ensure our algorithm is efficient, we sequentially run the exponential mechanism $k$ times to estimate and choose seeds similar to the greedy algorithm of Eckles et al. (2022), reducing their time complexity to $O(nmk)$. LABEL:thm:exp:inf shows that our exponential mechanism for single seed selection, $\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}},S)$, is $\epsilon$-ISDP (Definition 2.4), and provides a tail bound on the approximation error of maximizing $I_{\mathbf{x}}$. Its proof in Section 7.3 follows from the facts that $I_{\mathbf{x}}$ is a submodular function, and the exponential mechanism in Algorithm 3 can be seen as a noisy maximum function which implements an approximate greedy step by LABEL:thm:exp:utility.
Theorem 3.1 (label = thm:exp:inf)
Given $k\leq n\in\mathbb{N}$, $\epsilon>0$ and $m\in\mathbb{N}$ influence samples, ${\mathbf{x}}$, $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ is $\epsilon$-ISDP and outputs a set of $k$ nodes which, for all $t>0$, satisfies:
$\Pr\left[I_{\mathbf{x}}(\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}}))\leq(1-e^{-1})\max_{S:|S|=k}I_{\mathbf{x}}(S)-\frac{2k^{2}n}{\epsilon m}(\ln n+t)\right]\leq ke^{-t},$
where the probability is over the randomness of the algorithm run for fixed ${\mathbf{x}}$.
The above theorem provides a tail bound on the approximation error of $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ with respect to the optimal value on $I_{\mathbf{x}}$, given a fixed influence sample ${\mathbf{x}}$. Equipped with Proposition 2.1, we can derive the following error bound with respect to the optimal value $I_{\mathcal{G}}$ of the influence maximization problem (proved in Section 7.4):
Theorem 3.2 (label = thm:exp:acc)
Given graph ${\mathcal{G}}$ on $n\geq 2$ nodes and $k\leq n$, for any $0<\alpha\leq 1$ and $\epsilon>0$, $\mathcal{M}^{k,\epsilon}_{exp}(\cdot)$ that runs Algorithm 1 but replaces the single seed selection steps of Algorithm 2 by $\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}},S)$, its exponential mechanism variant in Algorithm 3, is $\epsilon$-ISDP and its output infects at least $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes with high probability, using at least $m=\max\{\frac{12}{\alpha\epsilon},\frac{9}{\alpha^{2}}\}k\ln n$ influence samples in $O(knm)$ run time.
Note that if we are not concerned with privacy and set $\epsilon\gg\alpha$, then the number of samples to achieve $\alpha n$ additive error is of order $\alpha^{-2}k\ln n$, which improves upon the $k^{2}(\ln n/\alpha)/\alpha^{3}$ bound of Eckles et al. (2022). Moreover, our result does not violate Theorem 3 of Eckles et al. (2022) that shows $\Omega(k^{2})$ influence samples are required to achieve $(1-1/e)\mbox{\sf OPT}-\alpha n$ performance bound for all $\alpha<(1-1/e)/k$, because our algorithm also needs $\Omega(k^{2})$ samples when $\alpha<1/k$.
3.2 Local DP with Randomized Response
We can achieve a stronger notion of privacy, local DP, by adding noise to the input data before running our (or any other) algorithms. Line 3 of Algorithm 4 adds the requisite amount of noise to achieve local $\epsilon$-ISDP in the input dataset ${\mathbf{x}}$ — flipping each entry of ${\mathbf{x}}$ independently with probability $\rho_{\epsilon}=1/(1+e^{\epsilon})$ to generate the noisy input $\tilde{{\mathbf{x}}}$. However, the added input noise will bias our estimate of each candidate seed’s quality; in particular, $\operatorname*{\mathbb{E}}_{\tilde{{\mathbf{x}}}}[I_{\tilde{{\mathbf{x}}}}(S|T)]$ is not necessarily equal to $I_{\mathcal{G}}(S|T)$ for a pair of subsets $S,T\subset V$ and we no longer have the benefit of applying Proposition 2.1 directly. Rather in Algorithm 5 we propose a post-processing of the noisy input data $\tilde{{\mathbf{x}}}$ to derive an unbiased estimate of a seed set’s quality. Importance of proper post-processing steps in the analysis of DP data are highlighted in other applications, e.g., census DP data release, where ignoring them can lead to biased decision-making and sub-optimal outcomes (Abowd et al. 2019, Ashmead et al. 2019, Zhu et al. 2022, 2021). The added variability of the estimators due to the DP noise can also degrade our estimation performance (even if not biased), which is a second concern in local DP algorithm design and opens a venue to explore alternative post-processing methods with potentially superior bias-variance tradeoff. In the remainder of this section, we first introduce Algorithms 4 and 5 and then offer LABEL:thm:rr:acc for their privacy and utility guarantees, addressing both concerns about the bias and variance of decision making with local DP influence samples.
Algorithm 4 Local $\epsilon$-ISDP seeding algorithm, $\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}})$
1:$m$ influence samples ${\mathbf{x}}$, total number of seeds $k$, and privacy budget $\epsilon>0$.
2:Set $S_{0}=\emptyset$.
3:Flip any bits of ${\mathbf{x}}$ with probability $\rho_{\epsilon}=\frac{1}{e^{\epsilon}+1}$ and get $\tilde{{\mathbf{x}}}$.
4:for $i\in[k]$ do
[…]
6:    $S_{i}=S_{i-1}\cup v_{i}$.
7:end for
8:Return $S_{k}$
Given a set $S$, an influence sample $x\in\{0,1\}^{n}$, and its perturbed version $\tilde{x}$, consider two sets of nodes $S\cap x$ and $S\cap\tilde{x}$. By Proposition 2.1, the influence function $I_{\mathcal{G}}(S)$ can be estimated from the reachable set, $n\Pr[S\cap x\neq\emptyset]$. But to achieve local DP, we can only observe the perturbed, reachable set $S\cap\tilde{x}$. In other words, let $\tilde{z}_{S}:=|S\cap\tilde{x}|$ and ${z}_{S}:=|S\cap{x}|$ be two random variables. If we can estimate $\Pr[z_{S}=0]$ from our observed samples of $\tilde{z}_{S}$, then we can estimate the size of the reachable set from $I_{\mathcal{G}}(S)=n\Pr[z_{S}>0]$ — using Proposition 2.1 — and then run the standard greedy algorithm for seed selection based on the estimated values.
Now, we discuss how to estimate $\Pr[z_{S}=0]$ from samples of $\tilde{z}_{S}$. The main observation is that the probability mass function of $\tilde{z}_{S}$ is a linear function of the probability mass function of $z_{S}$. Note that $S\cap x$ is the set of elements in $S$ that is also in the influence sample $x$. Hence, $S\cap\tilde{x}$ can be partitioned into (1) the subset $(S\cap\tilde{x})\cap x$ of elements of $x$ that remain in $S$ after perturbation, and (2) the subset $(S\cap\tilde{x})\setminus x$ of the elements of $\tilde{x}$ that join $S$ after the perturbation. The size of the first subset is a binomial random variable $Bin(|S\cap x|,1-\rho)$, and the size of the second subset is $Bin(|S\setminus x|,\rho)$. Therefore, given $|S|=l$ for all $a$ and $b$, the probability of $\tilde{z}_{S}=b$ conditional on $z_{S}=a$ is $\Pr[\tilde{z}_{S}=a|z_{S}=b]=\Pr[Bin(b,1-\rho)+Bin(l-b,\rho)=a],$ for all $a,b=0,\dots,l$ where $Bin(b,1-\rho)$ and $Bin(l-b,\rho)$ are independent. Additionally, if we set matrix $C=C_{\rho,l}\in{\mathbb{R}}^{(l+1)\times(l+1)}$ be a likelihood function where $C(a,b)=C_{\rho,l}(a,b)=\Pr[\tilde{z}_{S}=a|z_{S}=b]$, then we have:
$\begin{pmatrix}\Pr[\tilde{z}_{S}=0]\\ \Pr[\tilde{z}_{S}=1]\\ \vdots\\ \Pr[\tilde{z}_{S}=l]\end{pmatrix}=C_{\rho,l}\begin{pmatrix}\Pr[{z}_{S}=0]\\ \Pr[{z}_{S}=1]\\ \vdots\\ \Pr[{z}_{S}=l]\end{pmatrix}.$ (10)
We can write out each entry explicitly as follows:
$C_{\rho,l}(a,b)=\sum_{\ell=\max(0,b-a)}^{\min(l-a,b)}\binom{b}{\ell}\binom{l-b}{a-b+\ell}\rho^{a-b+2\ell}(1-\rho)^{l-a+b-2\ell}.$ (11)
To see why, condition on $\ell$ of the $z_{S}=b$ nodes in $S\cap x$ being flipped, leaving $b-\ell$ nodes in $S\cap x\cap\tilde{x}$. To get $\tilde{z}_{S}=a$ nodes in $S\cap\tilde{x}$, the remaining $a-b+\ell$ nodes should come from perturbing the status of the $l-b$ nodes belonging to $S\setminus x$. Note that our indexes $a$ and $b$ start at 00 instead of $1$. In the rest of the section, we will adopt this convention.
If matrix $C_{\rho,l}$ is invertible, then we can have an unbiased estimation of $\Pr[z_{S}>0]$ using Algorithm 5, below. LABEL:prop:feasible, proved in Section 7.5, shows that $C_{\rho.l}$ is invertible for $\rho>1/2$ and thus completes the specification of our randomized response algorithm.
Algorithm 5 Unbiased estimate for number of reachable set $J_{m}(S)$
1:$m$ perturbed influence samples $\tilde{{\mathbf{x}}}=(\tilde{x}^{1},\ldots,\tilde{x}^{m})$, their perturbation probability $\rho$, and a candidate set $S\subset V$ with size $|S|=l$.
2:Compute $\tilde{{\mathbf{f}}}=(\tilde{f}_{0},\tilde{f}_{1},\dots,\tilde{f}_{l})\in\mathbb{R}^{l+1}$ the empirical distribution of $\tilde{z}_{S}=|S\cap\tilde{x}|$ so that $\tilde{f}_{a}=\frac{1}{m}\sum_{t=1}^{m}\mathbf{1}[|S\cap\tilde{x}^{t}|=a]$ for all $a=0,1,\dots,l$.
3:Solve for ${\mathbf{f}}=(f_{0},f_{1},\dots,f_{l})\in{\mathbb{R}}^{l+1}$ so that $\tilde{{\mathbf{f}}}=C_{\rho,l}{\mathbf{f}}$.
4:Return $J_{m}(S)=n(1-f_{0})$
Proposition 3.3 (label = prop:feasible, name = Feasibility of Unbiased Post-processing)
Given any positive integer $l\geq 1$ and $\rho\in[0,1/2)$, matrix $C_{\rho,l}$ is invertible.
Note that when $\rho=1/2$, $\tilde{{\mathbf{x}}}$ consists of uniform random bits, and we cannot recover $\Pr[z_{S}>0]$ from the perturbed influence samples, and LABEL:prop:feasible shows that we can recover $\Pr[z_{S}>0]$ for any $\rho<1/2$.
Theorem 3.4 (label = thm:rr)
Given $\epsilon>0$, $m$ influence samples on a graph with $n$ nodes and integer $k<n$, Algorithm 4 is locally $\epsilon$-ISDP, and its output satisfies: $I_{\mathcal{G}}(\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}}))>(1-1/e)\mbox{\sf OPT}-\delta,$ with probability $1-2n^{k}\exp\left(-{m\delta^{2}}/{(2k^{2}n^{2}V_{\rho_{\epsilon},k}^{2})}\right),$ where $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})$.
The above theorem, proved in Section 7.6, provides a tail bound on the approximation error of $\mathcal{M}_{loc}^{k,\epsilon}$ given a fixed number of influence samples. We can convert the above result to a high probability error bound by choosing $m=\Omega({k^{3}V_{\rho_{\epsilon},k}^{2}\ln n}/{\alpha^{2}})$ which for $k$ constant and $\epsilon\gg\alpha$ improves on the $1/\alpha^{3}$ sample complexity in LABEL:thm:original (complete proof in Section 7.7):
Theorem 3.5 (label = thm:rr:acc)
Given graph ${\mathcal{G}}$ with $n\geq 2$ nodes and $k\leq n$, for any $0<\alpha\leq 1$ and $\epsilon>0$, $\mathcal{M}_{loc}^{k,\epsilon}$ is locally $\epsilon$-ISDP and its output infects at least $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes with high probability, using at least $m=O({k^{3}V_{\rho_{\epsilon},k}^{2}\ln n}/{\alpha^{2}})$ influence samples in $O(nk^{4}+kn^{2}m)$ run time, where $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})$.
Note that $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})$ in LABEL:thm:rr and LABEL:thm:rr:acc can be simplified as follows: $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})=O((1/2-1/(e^{\epsilon}+1))^{-k^{2}})=O([2(e^{\epsilon}+1)/(e^{\epsilon}-1)]^{k^{2}})$. Therefore $V_{\rho_{\epsilon},k}=O(\epsilon^{-k^{2}})$ as $\epsilon\to 0$ and $V_{\rho_{\epsilon},k}=O(2^{k^{2}})$ as $\epsilon\to\infty$.
4 The Nature of Our Privacy Protections in Social Networks
From the perspective of privacy guarantees, exponential mechanism (central DP) only protects the input data against the algorithmic revelations of the output (seeded nodes), giving the algorithm itself unhindered access to user data. On the other hand, the randomized response mechanism (local DP) provides a stronger privacy notion that protects the input data directly against any (adversarial) misuse, regardless of – and including — their use for seeding. Our results address both central and local notions of privacy, providing performance (utility) and sample complexity guarantees for influence maximization algorithms whose outputs/inputs are randomized at a desired level to give central/local DP protection given our input data structure, i.e., influence samples.
In our framework, we explicitly establish a budget $(\epsilon)$ for privacy leaks through the observed network cascades. This budget is a quantifiable limit on how much privacy/security risk one is willing to tolerate across all queries, not just a single query of the data. The latter highlights yet another way, in which our formulation of privacy budget is distinct from the setting of node- and edge-DP for graph input data structures. While algorithmically tractable, practice-wise our notion of adjacency protects each data donor against revelations of their contributions to any given cascade. While a donor may be willing to donate their contagion history, there may be some cascades in that history that they feel uncomfortable about (revealing sensitive or compromising information). Or the donors may dynamically contribute their contagion data where previous donations are irrevocable. Our notion of influence sample differential privacy provides donors with plausible deniability at any fixed cascade. This way donors are incentivized to contribute their entire set of $m$ cascades because we can bound their privacy risks from information leakage at any specific cascade in the set of size $m$, either centrally or locally - depending on the data curation environment - using our Algorithms 3 and 4. As long as the number of sensitive cascades is much smaller than $m$, we believe this notion of adjacency provide meaningful privacy protections. This is reflected in the requisite randomization in our greedy exponential mechanism (Algorithm 3): to achieve $\epsilon-$DP, each node $v$ is chosen with probability proportional to $\exp\left(\frac{\epsilon m}{2n}I_{\mathbf{x}}(v|S)\right)$, where $I_{\mathbf{x}}(v|S)$ is the marginal benefit of adding $v$ to the current seed set $S$ and it is estimated from the input influence samples ${\mathbf{x}}$. As $m\to 0$, this “soft max” converges to a uniform random choice, completely ignoring the input influence samples ${\mathbf{x}}$. On the other hand, the randomized response mechanism achieves $\epsilon-$DP locally by flipping the status of nodes in each cascade independently at random with probability $1/(1+e^{\epsilon})$. While the randomized response noise level is determined independently of the sample size $m$, the latter scales up as $m=O({k^{3}\epsilon^{-2k^{2}}\ln n}/{\alpha^{2}})$ with the decreasing privacy budget $(\epsilon)$ to sustain algorithmic performance in the privacy-critical regime ($\epsilon\to 0$).
As a concrete case, consider the choice of community leaders to undergo training for a public health or education intervention. To identify and target community leaders for treatment, the interventionist can survey the social network connections among the population. For example, Banerjee et al. (2013) survey individuals to collect social network data along 12 dimensions, including whom they get medical advice from or give medical advice to, relative and non-relatives with whom they socialize, and people with whom they exchange goods. While the community leaders may consent to and be properly compensated for their time in treatment, their social contacts face a privacy risk in having one of their associates chosen. As such, people may refrain from contributing to social network data collection or refuse to answer survey questions, especially when association with the treated individuals can pose harm, such as stigma or retaliation. For example, to prevent school bullying a few students may be recruited into treatment programs — e.g., by encouraging them to take a public stance against conflict (Paluck et al. 2016) — if the treated individuals are chosen based on their social network positions, and even if they are protected in treatment, their friends and contacts can face retaliation outside the treatment. Our central privacy protections in Section 3.1 allows the network participants to make a credible claim that their network associates are recruited into treatment by random choice, irrespective of their associations or contribution to the collected network cascade data.
In other related contexts, the nature of collected data can itself be compromising, irrespectively of how individuals are selected into treatments; e.g., surveillance and control of sexually transmitted disease and HIV may entail surveys of sexual contacts or needle sharing networks among people who inject drugs (Fairley et al. 2022, Endo et al. 2022). Administering such privacy-critical data should entail local privacy protections, whereby individual data is randomized as they are collected before any intervention designs or allocation decisions. Our local privacy guarantees in Section 3.2 are suited for such cases where individual contagion data may be collected for intervention design, but needs to be protected locally at the collection phase.
In a similar vein, social media platforms rely on social network data and information from past diffusions to organize their ad campaign and improve the efficacy of targeted advertising. Ensuring that these network operations do not violate users’ privacy is critical to the sustainability of social media platforms and their ad economies. Minority groups, in particular, can face more privacy risks than others in a social network. For example, a few adversarial nodes can coordinate their actions to move a web crawler towards or away from certain parts of an online social network to compromise their privacy through a re-identification attack (Backstrom et al. 2007). More broadly, deployment of learning systems to improve user experience requires accessing and aggregating historical usage information (e.g., for emojis and web domains) to optimize resource overhead for computation and communication, while respecting the personal and sensitive nature of the collected data (Apple Differential Privacy Team 2017, Erlingsson 2014b).
5 Simulation Results
Figure 1: We show the performance (expected influence $I_{\mathcal{G}}(\cdot)$) of the outputs of our exponential mechanism and randomized response algorithms, as well as the non-private, influence sample seeding algorithm and the complete-information, deterministic greedy baseline over a synthetic network generated using Erdős-Rényi model with parameters $n=200$ and $p=0.15$. We evaluate the expected spread size at different privacy budgets ($\epsilon$), as the number of influence samples ($m$) increases. Each point on the plot is an average of $50\times 20=1000$ trials where we generate $50$ collections of $m$ influence samples, and run our mechanisms on each collection $20$ times. The $50$ realizations capture the randomness in the diffusion process (influence sampling) and for each set of influence samples, we perform $20$ iterations to average out the randomness of the algorithms. The error bars show the $95\%$ confidence intervals of $1000$ trials around the means. Finally, to evaluate the performance of our mechanisms, we use another $1000$ influence samples to estimate the expected size of the spread.
Simulations on a synthetic input graph.
To test the performance of our algorithms in a sanitized setup, we generate an input using the Erdős-Rényi random graph model on $200$ nodes with edge probability $0.15$ and run our randomized response and exponential mechanism algorithms with different privacy budgets ($\epsilon=0.1,0.5$, or $1$) and different numbers of influence samples ($m$ ranging from zero to $500$). The independent cascade probability is set to $0.03$. Figure 1 shows that the rate of increase with $m$ slows down with the decreasing privacy budget for both exponential mechanism and randomized response algorithm alike.
Simulations on a real-world dataset.
We use a real dataset representing email interactions between individuals in a large European research institution taken from Stanford Large Network Dataset Collection (Stanford Network Analysis Project 2007, Yin et al. 2017, Leskovec et al. 2007). The nodes represent the individuals and the edges if there is communication between two individuals. This network represents the interaction between $1005$ nodes (individuals) through $25571$ edges. The interactions are mutual, and the graph is undirected. In the simulation setup of Fig. 2, we consider $k=4$ and $k=8$ seeds, the number of influence samples ranges from zero to $1500$, $m\in\{0,100,250,500,750,1000,1250,1500\}$, and the independent cascade probabilities are set to $p=0.0155$ on every edges in ${\mathcal{G}}$ and zero otherwise. The $m=0$ case corresponds to no input information where the choice of $k=4,8$ seeds will be uniformly at random. With no input information at $m=0$, there are no privacy risks to consider.
Figure 2: We test the performance (expected influence $I_{\mathcal{G}}(\cdot)$) of our privacy-preserving seeding algorithms (exponential mechanism and randomized response) on a real-world dataset, at different privacy budgets ($\epsilon$), as number of influence samples ($m$) increases. The figure also shows the non-private, influence sample seeding and the complete-information, deterministic, greedy baselines for comparison. Each point on the plot is average of $50\times 10=500$ trials where we generate $50$ collections of $m$ influence samples, and run our mechanisms on each collection $10$ times. The $50$ realizations capture the randomness in the diffusion process (influence sampling) and for each set of influence samples, we perform $10$ iterations to average out the randomness of the algorithms (randomized response and exponential mechanism). The error bars show the $95\%$ confidence intervals of $500$ trials around the means. Finally, to evaluate the performance of our mechanisms, we use another $2000$ influence samples to estimate the expected size of the spread.
In both Figures 1 and 2, we see that the exponential mechanism outperforms the randomized response mechanism as the latter provides stronger (local) privacy protections. We also observe that low values of privacy budget $\epsilon$ lead to earlier saturation in lower quality expected spread sizes. In those cases, our maximum number of influence samples is not enough to counteract the random noise introduced by the privacy mechanisms, so the value of network information diminishes with the decreasing privacy budge — cf., Eckles et al. (2022, Section 4). Increasing the privacy budget alleviates this effect, as we see a faster increase in the expected spread size with increasing $m$ at higher values of $\epsilon$. In fact, the performance of both mechanisms approaches the original (“non-private”) influence sample greedy seeding algorithm (Algorithm 1), as $\epsilon\to\infty$. For instance, in the randomized response mechanism (Algorithm 4), as the privacy budget increases, the flip probability of the entries in the influence samples $x\in\{0,1\}^{n\times m}$ (Algorithm 4, line 3) decreases, and matrix $C_{\rho,l}$ in Equation 11 is close to the identity matrix, so the performance is similar to the original influence sample greedy seeding algorithm (Algorithm 1). For the exponential mechanism of Section 3.1, at each step the mechanism approximately chooses an optimal seed that has the largest expected marginal contribution to the spread, and a large privacy budget ($\epsilon\to\infty$) sharpens the probability of choosing an optimal seed against others. Figures 1 and 2 also show the deterministic greedy baseline with complete network information to be compared with the exponential mechanism and randomized response algorithms at different $\epsilon$ values with increasing $m$. As we gain better information about the network by increasing the number of influence samples $(m\to\infty)$ and the privacy budget $(\epsilon\to\infty)$, the performance of both algorithms will converge to the deterministic greedy algorithm with complete network information.
6 Discussions, managerial insights and future directions
Comparison between node-, edge-, and influence-sample DP.
In this paper, we propose a new notion of differential privacy using influence samples adjacency. This notion may be extended to various problems which have underlying network structure. Examples include viral content and rumor cascades on social media, polls in elections from voters who influence each other’s opinions, and medical history of past disease contagions on epidemiological networks.
Note that the input of an influence sample DP algorithm is different from the inputs of node- or edge-DP algorithms – capturing different notions of privacy loss that are not formally comparable; specifically, in influence sample DP an individual may decide whether or not to report their particular value for an influence sample but they cannot change the data generation process. Instead of entry-level adjacency, one may consider individual-level adjacency that two inputs are adjacent if they differ at one row (an individual can decide whether or not to report their entire contagion history). Under this new notion, our analysis for exponential mechanism in Section 3.1 still hold, but we need to strengthen privacy constraints from $\epsilon$ to $\epsilon/m$ where $m$ is the number of samples, because an individual’s change on the influence sample increases from $1$ to $m$. However, as the privacy constraints depend on the number of samples, a large number of samples can no longer ensure that our mechanisms’ performance converges to the original greedy algorithm with full knowledge of the graph. This is inevitable as the new notion prevents our mechanisms from finding high-quality seeds regardless of the number of samples. Indeed, it is impossible to get any meaningful utility for ISDP with individual-level adjacency for similar reasons as edge-DP or node-DP (LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg).
Adjacency relation and differential privacy.
In this paper, we see a separation of utility guarantee between node/edge-DP algorithms and influence sample DP algorithms (LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg vs LABEL:thm:exp:acc and LABEL:thm:rr:acc). Such separations may be attributed to different notions of adjacency (node, edge, or influence sample adjacency). This observation suggests that differentially private algorithms can have significantly different utility and privacy guarantees by changing the notion of adjacency. However, we are not advocating influence sample DP as superior to node- or edge-DP. On the contrary, one goal of this paper is to bring attention to the role that different notions of adjacency play in discussions of privacy. For instance, because people are connected by social networks, their private data of any forms (census, medical records, etc.) are inevitably affected by the network structure. Those supporting node- and edge-DP may consider that the proper notion of adjacency for two datasets (on census, medical records, etc.) should also depend on the network rather than the traditional Hamming distance. Further discussion can be found in Rezaei and Gao (2019).
Efficiency and Robustness.
Our algorithms have polynomial running times in the size of the graph, with a time complexity of $O(knm)$ and $O(nk^{4}+kn^{2}m)$ for $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ and $\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}})$, respectively, for any constant privacy budget $\epsilon>0$. To improve running times, one may consider improving the iterative sampler for the exponential mechanism in Algorithm 3, or designing a better post-processing step for Line 3 in Algorithm 5. On the other hand, our ISDP randomized response algorithm $\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}})$ is automatically robust against independent randomness noise on influence samples. However, it is unclear how well our ISDP exponential mechanism $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ performs in the presence of noise.
Conclusions.
The correlated nature of social network information has been noted as a vulnerability in DP formulations which provide user-level guarantees with respect to the addition or removal of a single user’s data but ignore information leakage across different users’ records. By focusing on the utility of influence samples for influence maximization, our approach guarantees user’s participation in any given cascade is kept private either locally or centrally and preserves the statistical validity of the cascades for influence estimation and maximization. Subsequently, we provide a rigorous operationalization of action-level privacy by providing users with plausible deniability as to whether or not they have taken part in any particular network cascade (Kifer et al. 2020). This relaxation of user-level DP is ideally suited for the purposes of data collection to inform network intervention design. We expect that formulation and investigation of influence sample DP, with specific attention to the utility of the collected data for network intervention design, should lead to novel generalizations of privacy that are both consequentialist and robust to variations in sampling distributions and randomization noise (Kifer and Machanavajjhala 2014, Rastogi et al. 2009, Liu et al. 2020, Chakrabarti et al. 2022).
[…]
Rahimian and Yu contributed equally and are listed alphabetically. Rahimian is partially supported by the National Science Foundation under Grant No. 2318844. Rahimian also acknowledges support from a Pitt Momentum Funds Grant on Socially Responsible Data Collection and Network Intervention Designs. This research was supported in part by the University of Pittsburgh Center for Research Computing, RRID:SCR_022735, through the resources provided. Specifically, this work used the H2P cluster, which is supported by NSF award number OAC-2117681. Code for reported simulations can be accessed from https://github.com/aminrahimian/dp-inf-max

### v3

@@
 Seeding with Differentially Private Network Information
+[…]
 […]
@@
 […]
-When designing interventions in public health, development, and education, decision makers rely on social network data to target a small number of people, capitalizing on peer effects and social contagion to bring about the most welfare benefits to the population.
-Developing new methods that are privacy-preserving for network data collection and targeted interventions is critical for designing sustainable public health and development interventions on social networks.
-In a similar vein, social media platforms rely on network data and information from past diffusions to organize their ad campaign and improve the efficacy of targeted advertising.
-Ensuring that these network operations do not violate users’ privacy is critical to the sustainability of social media platforms and their ad economies.
-We study privacy guarantees for influence maximization algorithms when the social network is unknown, and the inputs are samples of prior influence cascades that are collected at random.
-Building on recent results that address seeding with costly network information, our privacy-preserving algorithms introduce randomization in the collected data or the algorithm output, and can bound each node’s (or group of nodes’) privacy loss in deciding whether or not their data should be included in the algorithm input.
-We provide theoretical guarantees of the seeding performance with a limited sample size subject to differential privacy budgets in both central and local privacy regimes.
-Simulations on synthetic and empirical network datasets reveal the diminishing value of network information with decreasing privacy budget in both regimes.
+In public health interventions such as the distribution of preexposure prophylaxis (PrEP) for HIV prevention, decision makers rely on seeding algorithms to identify key individuals who can amplify the impact of their interventions.
+In such cases, building a complete sexual activity network is often infeasible due to privacy concerns.
+Instead, contact tracing can provide influence samples, that is, sequences of sexual contacts without requiring complete network information.
+This presents two challenges: protecting individual privacy in contact data and adapting seeding algorithms to work effectively with incomplete network information.
+To solve these two problems, we study privacy guarantees for influence maximization algorithms when the social network is unknown and the inputs are samples of prior influence cascades that are collected at random and need privacy protection.
+Building on recent results that address seeding with costly network information, our privacy-preserving algorithms introduce randomization in the collected data or the algorithm output, and can bound the privacy loss of each node (or group of nodes) in deciding to include their data in the algorithm input.
+We provide theoretical guarantees of seeding performance with a limited sample size subject to differential privacy budgets in both central and local privacy regimes.
+Simulations on synthetic random graphs and empirically grounded sexual contacts of men who have sex with men reveal the diminishing value of network information with decreasing privacy budget in both regimes and graceful decrease in performance with decreasing privacy budget in the central regime.
+Achieving good performance with local privacy guarantees requires relatively higher privacy budgets that confirm our theoretical expectations.
 […]
-social network interventions, differential privacy, influence maximization, sample complexity, approximation algorithms, peer influence and social contagion, public health interventions
+social networks, differential privacy, influence maximization, sample complexity, approximation algorithms, sexual behavior, HIV, public health
 1 Introduction
-When designing interventions in public health, development, and education decision-makers rely on social network data to target a small number of people, capitalizing on peer effects and social contagion to bring about the most welfare benefits to the population.
-Many of these seeding methods are implemented without complete knowledge of social network structures — e.g., by targeting nominated friends of random individuals (Kim et al.
+When designing interventions in public health, development and education, decision makers rely on data from social networks to target a small number of people, leveraging peer effects and social contagion to deliver the greatest welfare benefits to the population.
+Many of these seeding methods are implemented without complete knowledge of social network structures – for example, by targeting nominated friends of random individuals (Kim et al.
 2015) or random pairs of friends (Alexander et al.
-2022) — and they enhance adoption of public health interventions by driving indirect treatment effects among people who were affected by spillovers from others’ treatment (even though they were not part of the initial treatment group).
-However, even without complete network information, network participants (and in particular minority members) who contribute to partial social network data collection endure privacy risks.
-This is especially important when engaging with vulnerable populations to record sensitive or stigmatizing behavior, e.g., in design of public health interventions among the homeless youth for HIV or suicide prevention (Wilder et al.
+2022) – and enhance adoption of public health interventions by driving indirect treatment effects among people who were affected by spillovers from others’ treatment (even though they were not part of the initial treatment group).
+However, even without complete network information, network participants (and particularly minority members) who contribute to partial social network data collection face privacy risks.
+This is especially important when vulnerable populations are involved in recording sensitive or stigmatizing behavior, such as the design of public health interventions among homeless youth for HIV or suicide prevention (Wilder et al.
 2021).
-Developing new methods that are privacy-preserving for network data collection and targeted interventions is critical for responsible design of public health interventions on social networks and maintaining community engagements that sustain such interventions.
-In this paper we propose randomization techniques based on the differential privacy formalism to provide social network participants with plausible deniability.
-Under these schemes the network participants can make a credible claim that their associates are targeted (e.g., recruited into treatment) by random choice, irrespective of their associations.
+In the context of HIV prevention, the distribution of pre-exposure prophylaxis (PrEP) can benefit from seeding algorithms that identify high-risk individuals based on data from social networks (Ross et al.
+2016, Simpson and Gumel 2017).
+However, constructing a complete sexual activity network for this purpose is challenging due to the sensitivity of the sexual behavior data.
+Instead, contact tracing can provide influence samples, which capture sequences of sexual activity without requiring full network knowledge.
+This introduces two key challenges: protecting the privacy of individuals contributing to contract tracing data (i.e., influence samples) and adapting seeding algorithms to effectively leverage these partial data.
+Developing privacy-preserving methods for network data collection and targeted interventions is critical for responsible public health interventions in social networks, particularly in vulnerable communities such as homeless youth, who are estimated to have a six to twelve times higher risk of HIV than housed youth (Caccamo et al.
+2017, Santa Maria et al.
+2020).
+Such methods ensure that privacy concerns do not undermine the trust and community engagement necessary for successful public health initiatives.
+In this paper, we propose randomization techniques based on the differential privacy formalism to provide social network participants with plausible deniability.
+Under these schemes, network participants can make a credible claim that their associates are targeted (e.g., recruited into treatment) by random choice, irrespective of their associations.
 We investigate how the targeting efficacy is affected by decreasing the privacy budget that entails more noise in the collected data (with local privacy mechanisms) or the targeting algorithm’s output (with central privacy mechanisms).
-After summarizing our main contributions in Section 1.1 and providing detailed comparisons with related works in Section 1.2, in Section 2 we provide preliminaries including hardness results that demonstrate the inadequacy of mainstream approaches to social network data privacy for seeding interventions and motivate our proposed notion of “influence sample differential privacy”.
-In Section 3, we provide privacy mechanisms for seeding in both central (Section 3.1) and local (Section 3.2) regimes.
+After summarizing our main contributions in Section 1.1 and providing detailed comparisons with related works in Section 1.2, in Section 2 we provide preliminaries including hardness results that demonstrate the inadequacy of mainstream approaches to the privacy of social network data for seeding interventions and motivate our proposed notion of “influence sample differential privacy”.
+In Section 3, we provide privacy mechanisms for seeding in central (Section 3.1) and local (Section 3.2) regimes.
 In Section 4, we discuss and contextualize the nature of our privacy protections for social network data collection and intervention design applications.
-In Section 5, we test the performance of our privacy mechanisms on synthetic and real social network datasets.
-We end with a discussion of limitations, future directions and concluding remarks in Section 6.
+In Section 5, we test the performance of our privacy mechanisms on synthetic Erdős-Rényi networks and empirically-grounded networks of men who have sex with men (MSM).
+We end with a discussion of limitations, future directions, and the concluding remarks in Section 6.
 Complete proofs and supplementary results are provided in several appendices at the end.
@@
 2014).
-In mathematical formulation, a randomized mechanism $\mathcal{M}$ is $\epsilon$-DP if for any two adjacent input datasets $D$ and $D^{\prime}$ in its input domain and any subset $A$ in its output range we have:
+In mathematical formulation, a randomized mechanism $\mathcal{M}$ is $\epsilon$-DP if for any two adjacent input datasets $D$ and $D^{\prime}$ in the input domain and any subset $A$ in the output range we have:
 $\frac{\Pr[\mathcal{M}(D)\in A]}{\Pr[\mathcal{M}(D^{\prime})\in A]}\leq e^{\epsilon}.$ (1)
-The $\epsilon$-DP criterion in Equation 1 limits inferences of any observer/adversary about the input data based on the algorithm output.
-The adjacency notion between datasets is typically defined such that neighboring datasets differ in contributions of an individual entity, and thus $\epsilon$-DP protects individuals’ contributions from ramification of an adversarial inference.
-Hence, different notions of adjacency will imply different protections for the individuals.
-In the sequel, we explain our proposed notion of adjacency for seeding with randomized contagion data (which we term “influence samples”), and the nature of privacy protections that it offers.
-Throughout this paper we use DP to signify any of the phrases “differentially private”, “differentially privately”, and “differential privacy” with or without hyphen between the words as appropriate by the language function that these words play in the context that DP is used.
+The $\epsilon$-DP criterion in Equation 1 limits the inferences of any observer/adversary about the input data based on the algorithm output.
+The notion of adjacency between datasets is typically defined so that neighboring datasets differ in contributions of an individual entity, and thus $\epsilon$-DP protects individuals’ contributions from ramification of an adversarial inference.
+Hence, different notions of adjacency will imply different protections for individuals.
+In the sequel, we explain our proposed notion of adjacency for seeding with randomized contagion data (which we term “influence samples”) and the nature of privacy protections that it offers.
+Throughout this paper, we use DP to signify any of the phrases “differentially private”, “differentially privately”, and “differential privacy” with or without hyphen between the words as appropriate by the language function that these words play in the context that DP is used.
 Seeding with partial and private network information.
-We study a classical, NP-hard network optimization problem, known as “influence maximization”, which asks to select $k$ nodes on a graph of size $n$ such that the expected spread size from seeding those $k$ nodes is maximum under a randomized model of network diffusion, whereby each edge passes on the influence (or “infection”) with an independent cascade probability $p$.
+We study a classical, NP-hard network optimization problem, known as “influence maximization”, which asks to select $k$ nodes on a graph of size $n$ such that the expected spread size from seeding those $k$ nodes is maximum in a randomized network diffusion model, whereby each edge passes on the influence (or “infection”) with an independent cascade probability $p$.
 Given the social network graph ${\mathcal{G}}$ and an integer $k\geq 1$, we denote the expected spread size from infecting the nodes in $S$ by $I_{\mathcal{G}}(S)$ and write $\mbox{\sf OPT}=\max_{S:|S|\leq k}I_{\mathcal{G}}(S)$ for the optimum solution to the $k$-influence maximization problem.
 It is an example of a cardinality-constrained combinatorial optimization on a graph with a monotone, submodular objective that admits a tight $(1-1/e)$ approximation guarantee using the greedy node selection algorithm.
-This problem has occupied a central place in graph and social networks algorithms literature since its introduction by Kempe et al.
+This problem has occupied a central place in the graph and social networks algorithms literature since its introduction by Kempe et al.
 (2003).
@@
 However, the graph structure of the input data has stunned the design of DP influence maximization algorithms.
-Common operationalizations of DP for graph algorithms, such as node- or edge-DP define adjacent inputs in Equation 1 based on neighboring graphs differing in a node or an edge.
-This adjacency criteria is often too stringent for graph problems whose solutions are extremely sensitive to existence of nodes and edges (see Section 1.2 for detailed citations and comparisons).
+Common operationalizations of DP for graph algorithms, such as node or edge DP, define adjacent inputs in Equation 1 based on neighboring graphs differing in a node or an edge.
+This adjacency criterion is often too stringent for graph problems whose solutions are extremely sensitive to the existence of nodes and edges (see Section 1.2 for detailed citations and comparisons).
 Our main hardness results indicate that it is impossible to obtain useful performance guarantees for influence maximization subject to node- or edge-DP constraints (see Section 2.2, LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg, summarized in 1 below).
 Informal Main Results 1
-Given $n\geq 1$ and $\epsilon>0$, there can be no $\epsilon$-node-DP or $\epsilon$-edge-DP seeding algorithms that give an approximation guarantee better than $(1-1/e)\mbox{\sf OPT}-\alpha_{\epsilon}{n}$ for $\alpha_{\epsilon}\geq\frac{1}{1000e^{\epsilon}}$.
+Given $n\geq 1$ and $\epsilon>0$, there can be no $\epsilon$ node-DP or $\epsilon$ edge-DP seeding algorithms that give an approximation guarantee better than $(1-1/e)\mbox{\sf OPT}-\alpha_{\epsilon}{n}$ for $\alpha_{\epsilon}\geq\frac{1}{1000e^{\epsilon}}$.
 A new privacy notion for social network data collection and intervention design.
 Our main contribution here is to formulate a new differential privacy notion with guarantees for influence maximization when the social network is unknown and the algorithm inputs are samples of prior influence cascades that are collected at random, i.e., “influence samples” — rather than the entire social network graph.
-Our results build on recent work about seeding with costly network information (Eckles et al.
-2022), and more broadly sample complexity of influence maximization (Sadeh et al.
+Our results are based on recent work on seeding with costly network information (Eckles et al.
+2022), and more generally sample complexity of influence maximization (Sadeh et al.
 2020, Borgs et al.
 2014).
-Our algorithm input is a set of $m$ influence samples, i.e., $m$ independent cascade samples that are collected on the social network graph each starting from a random, initially-infected node.
-We denote this input deadset by a binary matrix ${\mathbf{x}}\in\{0,1\}^{n\times m}$, where each row represents a node and each column represents the status of the node in an influence sample: $[{\mathbf{x}}]_{ij}=1$ if node $i$ is infected in the $j$-th influence sample and $[{\mathbf{x}}]_{ij}=0$ otherwise.
-Our privacy notion is with respect to the contribution of individual network participants to the influence samples, hence, we define the adjacent input datasets in Equation 1 as ${\mathbf{x}},{\mathbf{x}}^{\prime}\in\{0,1\}^{n\times m}$ which differ in a single binary entry:
+Our algorithm input is a set of $m$ influence samples, i.e., $m$ independent cascade samples that are collected on the social network graph, each starting from a random, initially infected node.
+We denote this input data set by a binary matrix ${\mathbf{x}}\in\{0,1\}^{n\times m}$, where each row represents a node and each column represents the status of the node in an influence sample: $[{\mathbf{x}}]_{ij}=1$ if node $i$ is infected in the $j$-th influence sample and $[{\mathbf{x}}]_{ij}=0$ otherwise.
+Our notion of privacy is with respect to the contribution of individual network participants to the influence samples; hence, we define the adjacent input data sets in Equation 1 as ${\mathbf{x}},{\mathbf{x}}^{\prime}\in\{0,1\}^{n\times m}$, which differ in a single binary entry:
 Influence Sample Differential Privacy (ISDP)We define “$\epsilon$ influence sample differential privacy ($\epsilon$-ISDP)” by controlling the odds ratio of the output of the seeding algorithm on adjacent influence samples, i.e., those differing in participation of a single node in a single influence cascade.
 We follow the formulation of influence maximization by Eckles et al.
-(2022) where the social network is unknown and influential nodes need to be estimated from costly samples of past spreads.
+(2022) where the social network is unknown and the influential nodes must be estimated from costly samples of past spreads.
 Our privacy guarantees are with respect to influence samples, controlling information leakage about the cascade data.
 For example, Apple Inc.
-uses a randomized response mechanism to collect the users’ daily usage of emojis with a fixed privacy budget ($\epsilon=4$).
+uses a randomized response mechanism to collect the users’ daily usage of emojis with a fixed privacy budget ($\epsilon=4$); see Apple Differential Privacy Team (2017).
 This setup maps naturally to our formulation of ISDP as the usage of each emoji can be modeled as an influence sample.
@@
 Sample complexity with differential privacy and approximation guarantees.
-By introducing randomization at the input or output of seeding algorithms, we are able to bound each node’s privacy loss in deciding whether or not their data should be included in the influence samples — as opposed to the social network graph.
-Of note, this is critically different from DP analysis of graphs that is defined with respect to neighboring graphs in the algorithm input.
-In our approach, each node’s decision to appear in the input does not affect the data generation process (DGP) of the input sample or the underlying diffusion process; the network graph remains the same and the privacy implications are studied with respect to the construction of the samples on a fixed network.
+By introducing randomization at the input or output of seeding algorithms, we are able to bound each node’s privacy loss in deciding whether or not their data should be included in the influence samples, as opposed to the social network graph.
+Notably, this is critically different from DP analysis of graphs that is defined with respect to neighboring graphs in the algorithm input.
+In our approach, each node’s decision to appear in the input does not affect the data generation process (DGP) of the input sample or the underlying diffusion process; the network graph remains the same, and the privacy implications are studied with respect to the construction of the samples on a fixed network.
 This allows us to meaningfully bound the effect of the removal of a node’s data on the algorithmic performance and to use those bounds to tradeoff the privacy risks against improved performance.
-To mitigate privacy risks, DP uses randomization to provide plausible deniability certificates to those contributing to the input data.
-This is achieved in one of the two ways: (i) using exponential mechanism (McSherry and Talwar 2007) by randomizing the algorithm output, i.e., central DP, (ii) using the randomized response mechanism (Warner 1965) by injecting noise in the input data, i.e., local DP.
-Accordingly, we propose two efficient DP seeding algorithms and bound their performance in terms of the sample size $m$, network size $n$, seed set size $k$, and privacy budget $\epsilon$, (See Section 3 LABEL:thm:exp:acc and LABEL:thm:rr:acc):
+To mitigate privacy risks, DP uses randomization to provide plausible deniability certificates to those who contribute to the input data.
+This is achieved in one of the two ways: (i) using the exponential mechanism (McSherry and Talwar 2007) by randomizing the algorithm output, i.e., central DP; (ii) using the randomized response mechanism (Warner 1965) by injecting noise in the input data, i.e., local DP.
+Consequently, we propose two efficient DP seeding algorithms and bound their performance in terms of sample size $m$, network size $n$, seed set size $k$, and privacy budget $\epsilon$, (see Section 3 LABEL:thm:exp:acc and LABEL:thm:rr:acc):
 Informal Main Results 2
@@
 Compared to our central DP seeding algorithm, the local algorithm requires more (exponentially in $k$) influence samples: $m=O({k^{3}V_{\rho_{\epsilon},k}^{2}\ln n}/{\alpha^{2}})$ with $O(nk^{4}+kn^{2}m)$ run time, where $V_{\rho_{\epsilon},k}=O(\max\{2,1/\epsilon\}^{k^{2}})$.
-Given $m$ influence samples, in Section 3.1 we use exponential mechanisms (McSherry and Talwar 2007) to choose high quality seeds with central DP guarantees.
-In Subsection 3.2, before accessing the data (influence samples), we use the randomized response mechanism (Warner 1965), to perturb the input data and protect privacy at the local level.
+Given $m$ influence samples, in Section 3.1 we use exponential mechanisms (McSherry and Talwar 2007) to choose high-quality seeds with central DP guarantees.
+In Section 3.2, before accessing the data (influence samples), we use the randomized response mechanism (Warner 1965), to perturb the input data and protect privacy at the local level.
 However, these perturbations can introduce additional biased noise into the data, so we provide a closed-form characterization of the introduced noise and an algorithm to offset this bias.
-In Section 5, we thoroughly evaluate our algorithms on synthetic and real graphs of different sizes and verify our theoretical guarantees in simulations.
+In Section 5, we thoroughly evaluate our algorithms on real and synthetic graphs of different sizes and verify our theoretical guarantees in simulations.
 1.2 Comparisons with related work
@@
 A long line of literature, going back to the seminal works of Kempe et al.
-(2003, 2005, 2015) propose efficient approximation algorithms for this problem based on the submodularity of the influence function (Chen et al.
+(2003, 2005, 2015) proposes efficient approximation algorithms for this problem based on the submodularity of the influence function (Chen et al.
 2009, Wang et al.
 2012).
-This problem arise naturally in applications such as development economics and public health to facilitate adoption of welfare-enhancing technologies, e.g., microfinance (Banerjee et al.
+This problem arises naturally in applications such as development economics and public health to facilitate the adoption of welfare-enhancing technologies, e.g., microfinance (Banerjee et al.
 2013), insurance (Cai et al.
 2015), contraception (Behrman et al.
-2002), water purification methods, and multivitamins (Kim et al.
-2015); and to target localized social network interventions for spreading information about immunization camps (Banerjee et al.
-2019), preventing misinformation about drug side effects (Chami et al.
-2017), increasing HIV awareness among homeless youth (Yadav et al.
-2017), and reducing bullying among adolescent youth (Paluck et al.
+2002), water purification methods and multivitamins (Kim et al.
+2015); and to target localized social network interventions to spread information about immunization camps (Banerjee et al.
+2019), prevent misinformation about drug side effects (Chami et al.
+2017), increase awareness of HIV among homeless youth (Yadav et al.
+2017, 2018), and reduce bullying among adolescents (Paluck et al.
 2016).
@@
 Kim et al.
-(2015) show similar results for a cluster-randomized trial in villages in Honduras, where they measure the efficacy of social network targeting strategies to promote adoption of chlorine for water purification and multivitamins for micronutrient deficiencies.
+(2015) show similar results for a cluster-randomized trial in villages in Honduras, where they measure the effectiveness of social network targeting strategies to promote the adoption of chlorine for water purification and multivitamins for micronutrient deficiencies.
 In all such settings, collecting complete social network connection (edge) data is difficult, costly, or impossible and comes with critical privacy and security risks to individuals in the social network.
-In this paper we focus on differential privacy guarantees for influence maximization when social network is unknown and influential nodes need to be estimated from costly samples of past spreads.
+In this paper, we focus on differential privacy guarantees for influence maximization when social network is unknown and influential nodes need to be estimated from costly samples of past spreads.
 Social network data privacy.
-Much of the past work on social network data privacy focuses on identifiability of nodes from released graph data, with or without additional information on node attributes and their neighborhoods (Romanini et al.
+Much of the previous work on social network data privacy focuses on the identifiability of nodes from released graph data, with or without additional information on node attributes and their neighborhoods (Romanini et al.
 2021).
-Classical anonymization techniques mask node attributes and perturb, modify, randomize or aggregate the graph structure to prevent re-identification of nodes within a confidence level — typically achieving k-anonymity (Wu et al.
+Classical anonymization techniques mask the attributes of the nodes and perturb, modify, randomize, or aggregate the structure of the graph to prevent reidentification of the nodes within a confidence level, typically achieving $k$ anonymity (Wu et al.
 2010, Zheleva and Getoor 2011, Abawajy et al.
@@
 However, statistical disclosure control of collected data requires safeguarding data donor’s privacy against adversarial attacks where common anonymization techniques are shown to be vulnerable to various kinds of identification (Barbaro et al.
-2006), linkage and cross-referencing (Narayanan and Shmatikov 2008, Sweeney 2015, 1997), statistical difference and re-identification attacks (Kumar et al.
+2006), linkage and cross-referencing (Narayanan and Shmatikov 2008, Sweeney 2015, 1997), statistical difference and reidentification attacks (Kumar et al.
 2007).
-On the other hand while edge sampling at the data collection can somewhat mitigate the re-identification risk of nodes (Romanini et al.
+However, while edge sampling at data collection can somewhat mitigate the risk of reidentification of nodes (Romanini et al.
 2021, Jiang et al.
-2021), the seeded nodes can reveal additional side information about who has contributed to network data collection.
-For example, once a community leader is chosen, his or her close contacts may suffer privacy leaks, even though no social network data are published.
+2021), seeded nodes can reveal additional side information about who has contributed to data collection.
+For example, once a community leader is chosen, his or her close contacts may suffer privacy leaks even though no social network data is published.
 Our privacy guarantees for influence maximization follow the DP formalism to provide network participants and data donors with plausible deniability with regard to their contribution to the collected data or their role in shaping the algorithm output (the seeded nodes).
@@
 DP analysis of graphs.
-Differential privacy (DP) is a versatile framework for developing privacy-preserving algorithms, and it has been impactful in large-scale and high-profile applications such as the U.S.
-decennial census Garfinkel (2022) and the tech industry Erlingsson (2014a), Apple Differential Privacy Team (2017).
-It have been successfully applied to a broad variety of problem classes, including, e.g., combinatorial (Gupta et al.
+Differential privacy (DP) is a versatile framework for developing privacy-preserving algorithms and has been influential in large-scale and high-profile applications such as the US Census Garfinkel (2022) and the tech industry Erlingsson (2014a), Apple Differential Privacy Team (2017).
+It has been successfully applied to a broad variety of problem classes, including, e.g., shortest path (Fioretto et al.
+2024), combinatorial (Gupta et al.
 2010), submodular (Mitrovic et al.
 2017, Sadeghi and Fazel 2021), and online submodular (Salazar and Cummings 2021) optimization problems.
-The existing literature on DP analyses of graphs, e.g., edge-DP Nissim et al.
-(2007) or node-DP (Kasiviswanathan et al.
-2013, Raskhodnikova and Smith 2016), has devoted much attention to graph statistics and parameters such as average degree and various subgraph counts (Blocki et al.
+The existing literature on graph DP analysis, for example, edge DP (Nissim et al.
+2007, Eden et al.
+2023) or node DP (Kasiviswanathan et al.
+2013, Raskhodnikova and Smith 2016, Kalemaj et al.
+2023), has devoted much attention to graph statistics and parameters such as average degree and various subgraph counts (Blocki et al.
 2022), degree distribution (Hay et al.
-2009) and the graph spectra (Wang et al.
-2013), as well as community Mohamed et al.
-(2022) and densest subgraph detection (Nguyen and Vullikanti 2021).
-These techniques are not directly applicable to influence maximization, and our main hardness results indicate that node- or edge-DP are too stringent to render useful performance guarantees for influence maximization.
-Similar observations have been made about node- and edge-DP guarantees of other graph algorithms which are too sensitive to the presence of individual edges (and especially nodes) in their input graphs (Gupta et al.
+2009) and graph spectra (Wang et al.
+2013), as well as community detection Mohamed et al.
+(2022) and densest subgraph detection problems (Nguyen and Vullikanti 2021).
+These techniques are not directly applicable to influence maximization, and our main hardness results indicate that node or edge DP is too stringent to render useful performance guarantees for influence maximization.
+Similar observations have been made about the node and edge DP guarantees of other graph algorithms which are too sensitive to the presence of individual edges (and especially nodes) in their input graphs (Gupta et al.
 2010, Li et al.
@@
 (2010) consider several graph problems including vertex and set cover.
-Their modified greedy algorithm for set cover is similar to our central DP mechanism in Section 3.1, though addressing a substantially different privacy environment, where the goal is to protect a private subset $R$ while outputting a minimal collection that covers $R$.
+Their modified greedy algorithm for set cover is similar to our central DP mechanism in Section 3.1, although it addresses a substantially different privacy environment, where the goal is to protect a private subset $R$ while outputting a minimal collection that covers $R$.
 In both cases direct application of the exponential mechanism is inefficient and the greedy approach facilitates an efficient implementation.
-In formulating our notion of “influence sample differential privacy” (ISDP), we move away from having the social network as the algorithm input.
+In formulating our notion of “influence sample differential privacy” (ISDP), we are moving away from having the social network as input to the algorithm.
 Instead, we focus on an input dataset of $m$ cascades that are collected at random on the social network graph.
 ISDP allows us to limit the information leakage about contribution of individual nodes to the input cascades (with local ISDP), or contribution of individual nodes to the algorithm output through the input cascades (with central ISDP).
-It is worth noting that achieving local privacy under ISDP is easy by randomizing individual statuses of the nodes in the input cascades.
-However, in general it is challenging to coordinate randomization across nodes in a meaningfully decentralized manner that matches the problem context, and locally deferentially private analysis of graphs has been limited so far to certain restricted settings (Imola et al.
+It should be noted that achieving local privacy under ISDP is easy by randomizing individual statuses of the nodes in the input cascades.
+However, in general, it is challenging to coordinate randomization between nodes in a meaningfully decentralized manner that matches the problem context, and the locally deferentially private analysis of graphs has been limited to certain restricted settings so far (Imola et al.
 2021, Dhulipala et al.
@@
 Privacy of network diffusions.
-The fundamental relationship between information flow and privacy is well noted, cf., contextual integrity (Nissenbaum 2009, Benthall and Cummings 2022).
-Notwithstanding, the privacy implications of information diffusion and algorithmic intervention on social networks is largely unexplored, with the exception of two studies (Rezaei et al.
+The fundamental relationship between information flow and privacy is well observed, cf.
+contextual integrity (Nissenbaum 2009, Benthall and Cummings 2024).
+However, the privacy implications of information diffusion and algorithmic intervention on social networks remain largely unexplored, with the exception of a few studies (Rezaei et al.
 2021, Liell-Cock et al.
+2020, Fanti et al.
+2017).
+The existing work looks at this issue in specific, highly stylized scenarios.
+Rezaei et al.
+(2021) show that it is extremely difficult to hide the existence of a giant connected component of infected nodes under an independent cascade model, and a simple inference attack can reveal the status of a good fraction of nodes (Rezaei et al.
+2021).
+Liell-Cock et al.
+(2020) propose a decaying pseudo-noise that is added locally to mask the influence structure against an external observer with access to the entire information flow under the Friedkin-Johnsen influence propagation model (Liell-Cock et al.
 2020).
-The existing work looks at this issue in specific, highly-stylized scenarios.
-One study shows that it is extremely hard to hide the existence of a giant, connected component of infected nodes under an independent cascade model and a simple inference attack can reveal the status of a good fraction of nodes (Rezaei et al.
-2021).
-In another study, the authors propose a decaying pseudo-noise that is added locally to mask the influence structure against an external observer with access to the entire information flow under the Friedkin-Johnsen influence propagation model (Liell-Cock et al.
-2020).
-With our formulation of DP for influence maximization we can focus on privacy leaks through information flow on the network rather than the underlying network structure.
-Our formulation of DP for influence samples ($\epsilon$-ISDP in 1) allows us to explicitly quantify information leaks from observation of random spreads on the network and for us to take a direct approach to the privacy risks that social network participants face from network diffusions.
+Motivated by the need to hide the identities of users who post sensitive messages, Fanti et al.
+(2017) propose a new messaging protocol called adaptive diffusion, which works by breaking the directional symmetry of diffusion around its true source and achieves fast spread and perfect source obfuscation for infinite regular trees, in the sense that all users with the message are equally likely to have been its originator.
+With our formulation of DP for influence maximization, we can focus on privacy leaks through the information flow in the network rather than the underlying network structure.
+Our formulation of DP for influence samples ($\epsilon$-ISDP in 1) allows us to explicitly quantify information leaks from the observation of random spreads on the network and to take a direct approach to the privacy risks that social network participants face from network diffusions.
 2 Problem Setup and Preliminary Results
@@
 The set of nodes, $V$, is indexed by $[n]:=\{1,\ldots,n\}$ and $p:V\times V\to[0,1]$ assigns probability weights to every edge in ${\mathcal{G}}$.
-An infection begins at a set $S$ of seed nodes and spreads through the graph in rounds.
-Each infected node, once they become infected, has a single chance of passing the infection to their neighbors.
-Each undirected edge $e=(v,u)$ has a weight $p_{e}\in[0,1]$ representing the probability that the process spreads along edge $e$ to node $u$ in the round immediately after $v$ is first infected.
+An infection begins in a set $S$ of seed nodes and spreads through the graph in rounds.
+Each infected node, once it becomes infected, has a single chance of passing the infection to their neighbors.
+Each undirected edge $e=(v,u)$ has a weight $p_{e}\in[0,1]$ that represents the probability that the process spreads along edge $e$ to node $u$ in the round immediately after $v$ is first infected.
+Some notation.
 The above process has the following equivalent formulation: we first sample an undirected, unweighted graph $g$ by independently realizing each edge $e$ with probability $p_{e}$.
 We write $g\sim{\mathcal{G}}$ to denote that $g$ is drawn from the random graph distribution ${\mathcal{G}}$.
-Then we can construct the set of infected nodes in the independent cascade process, as the set of nodes that are reachable from seed set $S$ in $g$.
-Now we define some notations.
+Then we can construct the set of infected nodes in the independent cascade process, as the set of nodes that can be reached from the seed set $S$ in $g$.
 Given an undirected graph $g$ and subsets of nodes $S$ and $T$, we write $C_{g}(S)$ for the set of nodes that are reachable from (or infected by) $S$ in $g$, and $C_{g}(S|T):=C_{g}(S)\setminus C_{g}(T)$ for the set of nodes that are reachable from $S$ but not from $T$.
-Then the influence function $I_{g}(S):=|C_{g}(S)|$ is the number of nodes reachable (or infected) by $S$ and $I_{\mathcal{G}}(S):=\operatorname*{\mathbb{E}}_{g\sim{\mathcal{G}}}[I_{g}(S)]$ is its expectation.
-Similarly, we define the conditional influence function $I_{g}(S|T):=|C_{g}(S|T)|$ and denote its expected value by $I_{\mathcal{G}}(S|T):=\operatorname*{\mathbb{E}}_{g\sim{\mathcal{G}}}[I_{g}(S|T)]$.
+The influence function $I_{g}(S):=|C_{g}(S)|$ is the number of nodes reachable (or infected) by $S$ and $I_{\mathcal{G}}(S):=\operatorname*{\mathbb{E}}_{g\sim{\mathcal{G}}}[I_{g}(S)]$ is its expectation.
+Similarly, we define conditional influence function $I_{g}(S|T):=|C_{g}(S|T)|$ and denote its expected value by $I_{\mathcal{G}}(S|T):=\operatorname*{\mathbb{E}}_{g\sim{\mathcal{G}}}[I_{g}(S|T)]$.
 The influence maximization problem.
@@
 Influence sample greedy seeding algorithm.
-How can we design a seeding algorithm $M$ that takes $m$ influence samples ${\mathbf{x}}$ and outputs a set of $k$ nodes, $M({\mathbf{x}})\in\mathcal{Y}:=\{0,1\}^{V}$?
+How can we design a seeding algorithm $M$ that takes $m$ influence samples ${\mathbf{x}}$ and produces a set of $k$ nodes, $M({\mathbf{x}})\in\mathcal{Y}:=\{0,1\}^{V}$?
 Here we present a seeding algorithm by Eckles et al.
 (2022).
-The core idea of Algorithm 1 is to estimate the influence function from the influence samples, and run the greedy algorithm.
+The core idea of Algorithm 1 is to estimate the influence function from the influence samples and to run the greedy algorithm.
 Specifically, given $m$ realized influence samples ${\mathbf{x}}$ and a subset of nodes $S$ in ${\mathcal{G}}$, we define the $S$-reachable influence samples as: $C_{\mathbf{x}}(S):=\{t\in[m]:\exists v\in S,x_{v}^{t}=1\},$ which is the set of influence samples whose activated nodes are reachable from the set $S$, and $C_{\mathbf{x}}(S|T)=C_{\mathbf{x}}(S)\setminus C_{\mathbf{x}}(T)$.
-Alternatively, if we treat ${\mathbf{x}}$ a collection of $m$ subset $x^{1},\dots,x^{m}$, $C_{\mathbf{x}}(S)=\{t:S\cap x^{t}\neq\emptyset\}$.
+Alternatively, if we treat ${\mathbf{x}}$ as a collection of $m$ subsets $x^{1},\dots,x^{m}$, $C_{\mathbf{x}}(S)=\{t:S\cap x^{t}\neq\emptyset\}$.
 Finally, we define $I_{\mathbf{x}}(S)=\frac{n}{m}|C_{\mathbf{x}}(S)|$ and $I_{\mathbf{x}}(S|T)=\frac{n}{m}|C_{\mathbf{x}}(S|T)|$.
-The following proposition says the expectation of $I_{\mathbf{x}}$ is equal to $I_{{\mathcal{G}}}$.
-We will use $v$ as $\{v\}$ for singleton sets to simplify the notation, e.g., $I_{\mathcal{G}}(v)$ as $I_{\mathcal{G}}(\{v\})$.
+The following proposition states that the expectation of $I_{\mathbf{x}}$ is equal to $I_{{\mathcal{G}}}$.
+We will use $v$ instead of $\{v\}$ for singleton sets to simplify the notation, for example, writing $I_{\mathcal{G}}(v)$ for $I_{\mathcal{G}}(\{v\})$.
 Proposition 2.1 (Observation 3.2. of Borgs et al. (2012) – cf. Borgs et al. (2014))
-Given a positive integer $m$, a graph ${\mathcal{G}}=(V,p)$ and sets of nodes $S,T\subseteq V$, if ${\mathbf{x}}$ is the random variable of $m$ influence sample, then $\operatorname*{\mathbb{E}}_{{\mathbf{x}}}[I_{{\mathbf{x}}}(S)]=I_{\mathcal{G}}(S)$, and $\operatorname*{\mathbb{E}}_{\mathbf{x}}[I_{\mathbf{x}}(S|T)]=I_{\mathcal{G}}(S|T)$.
+Given a positive integer $m$, a graph ${\mathcal{G}}=(V,p)$, and sets of nodes $S,T\subseteq V$, if ${\mathbf{x}}$ is the random variable that encodes $m$ influence samples, then $\operatorname*{\mathbb{E}}_{{\mathbf{x}}}[I_{{\mathbf{x}}}(S)]=I_{\mathcal{G}}(S)$, and $\operatorname*{\mathbb{E}}_{\mathbf{x}}[I_{\mathbf{x}}(S|T)]=I_{\mathcal{G}}(S|T)$.
 Additionally, by Hoeffding’s inequality, for all $\epsilon>0$ we have: $\Pr[|I_{{\mathbf{x}}}(S)-I_{\mathcal{G}}(S)|\geq\epsilon n]\leq 2\exp(-2\epsilon^{2}m),$ where the probability is over the randomness in the generative model of influence samples $x$.
 With the above observation, Eckles et al.
-(2022) design a seeding algorithm that only uses influence samples and show that if the number of influence sample is of order $k^{2}(\log n/\alpha)/\alpha^{3}$, then Algorithm 1 is optimal up to $\alpha n$ additive error.
-Informally, Algorithm 1 is a greedy algorithm that sequentially query Algorithm 2 for the seed that maximizes the marginal improvement.
-In this paper, we will follow similar idea that combine greedy algorithm and an estimation algorithm for marginal improvement.
+(2022) design a seeding algorithm that only uses influence samples and show that if the number of influence samples is of the order $k^{2}(\log n/\alpha)/\alpha^{3}$, then Algorithm 1 is optimal up to the additive error $\alpha n$.
+Informally, Algorithm 1 is a greedy algorithm that sequentially queries Algorithm 2 for the seed that maximizes marginal improvement.
+In this paper, we will follow a similar idea that combines a greedy algorithm and an estimation algorithm for marginal improvement.
 Algorithm 1 Influence sample greedy seeding algorithm
@@
 […]
-In Algorithm 1, we slightly modify the INF-SAMPLE$(m,k)$ algorithm to sequentially call Algorithm 2 for selecting the next seed.
-Instead of collecting fresh influence samples at each greedy step, Algorithm 1 calls Algorithm 2 to reuse the influence samples and have much better influence sample complexity as a result.
+In Algorithm 1, we slightly modify the INF-SAMPLE$(m,k)$ algorithm to sequentially call Algorithm 2 to select the next seed.
+Instead of collecting fresh influence samples at each greedy step, Algorithm 1 calls Algorithm 2 to reuse the influence samples and, as a result, to have a much better influence sample complexity.
 We formally show this improvement in LABEL:thm:exp:acc.
 Theorem 2.2 (label = thm:original, name = Theorem 2 of Eckles et al. (2022))
-Given ${\mathcal{G}}$ of $n\geq 2$ nodes and $k\leq n$, for any $0<\alpha\leq 1$, the output of Algorithm 1 can infect $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes using no more than $81k^{2}\log(6nk/\alpha)/\alpha^{3}$ influence samples.
-It is well known that even with the full knowledge of the graph ${\mathcal{G}}$ no efficient algorithm can perform better than $(1-1/e)\mbox{\sf OPT}$ unless P equals NP.
+Given the graph ${\mathcal{G}}$ on $n\geq 2$ nodes, we want to seed $k\leq n$ nodes.
+For any $0<\alpha\leq 1$, the output of Algorithm 1 can infect $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes using no more than $81k^{2}\log(6nk/\alpha)/\alpha^{3}$ influence samples.
+It is well known that even with complete knowledge of the graph ${\mathcal{G}}$, no efficient algorithm can perform better than $(1-1/e)\mbox{\sf OPT}$ unless P equals NP.
 Eckles et al.
@@
 Let $10<\alpha<$ 1 be any constant.
-For any algorithm with $o(n)$ samples, there exists a weighted graph ${\mathcal{G}}$ so that the output of the algorithm can only infect less than $\alpha\mbox{\sf OPT}$ nodes.
+For any algorithm with $o(n)$ samples, there exists a weighted graph ${\mathcal{G}}$ so that the output of the algorithm can only infect fewer than $\alpha\mbox{\sf OPT}$ nodes.
 2.2 Influence Sample Differential Privacy
-Intuitively, a differentially private (DP) algorithm ensures that the algorithm outputs on two adjacent inputs are close to each other, making them hard to distinguish.
-However, there are multiple alternative definitions of adjacency relation on graph-structured input data.
-One classical notion is node DP where two graphs are adjacent if they are different by one node’s connections.
-Another one is edge DP where two graphs are adjacent if they differ by one edge (Hay et al.
+Intuitively, a differentially private (DP) algorithm ensures that the algorithms output on two adjacent inputs are close to each other, making them difficult to distinguish.
+However, there are multiple alternative definitions of the adjacency relation in graph-structured input data.
+One classical notion is node DP, where two graphs are adjacent if they are different by the connections of one node.
+Another is edge DP, where two graphs are adjacent if they differ by one edge (Hay et al.
 2009).
-In this paper, we would like to focus on DP seeding algorithms that use influence samples but not raw graph data.
-Instead of raw graph, we study algorithms that only use historical cascades data — i.e., influence samples — where two collections of $m$ influences samples ${\mathbf{x}}$ and ${\mathbf{x}}^{\prime}$ in $\{0,1\}^{n\times m}$ are (influence sample) adjacent if we can change one entry to convert $\mathbf{x}$ to $\mathbf{x}^{\prime}$, and write ${\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}$ if they are adjacent.
-Given the above adjacency relation on influence samples, we can define if a stochastic seeding algorithm $M$ is DP.
+In this paper, we focus on developing differentially private (DP) seeding algorithms that utilize influence samples, rather than relying on raw graph data, which is often impractical to obtain in sensitive contexts such as public health interventions for HIV prevention.
+In such cases, contact tracing using surveys of sexual activity may be preferable to collecting a complete network due to privacy concerns.
+The highly contextualized and variable nature of contact data makes a reliable network representation not well defined, even if it is desirable.
+Therefore, we abstract away form the network representation and work directly with influence samples, which, in the case of contact tracing for HIV prevention, track the spread of sexual activity over time without requiring complete network information.
+We study algorithms that only use these influence samples, where two collections of $m$ influence samples, ${\mathbf{x}}$ and ${\mathbf{x}}^{\prime}$ in $\{0,1\}^{n\times m}$, are “influence sample” adjacent if one entry can be modified to transform ${\mathbf{x}}$ into ${\mathbf{x}}^{\prime}$.
+We denote this adjacency by ${\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}$.
+Given this adjacency relation on influence samples, we define a stochastic seeding algorithm $M$ to be DP if the algorithm’s output does not change significantly when applied to adjacent influence samples.
+In the HIV example, by focusing on influence samples derived from sexual behavior, our DP algorithms provide privacy guarantees for participants in sensitive public health settings while identifying high-risk individuals for targeted interventions.
 Definition 2.4 (ISDP)
 Given $\epsilon\geq 0$, $n,m\geq 1$, and $k\in[n]$, a function $M^{k}:\{0,1\}^{n\times m}\to\mathcal{Y}$ is $\epsilon$-influence sample differentially private ($\epsilon$-ISDP, in short) if for all outputs ${\mathbf{y}}\in\mathcal{Y}$ of $k$ seeds, and all pairs of adjacent datasets (collection of influence samples), $\mathbf{x}\underset{is}{\sim}\mathbf{x}^{\prime}$, we have: $\Pr[M^{k}(\mathbf{x})={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}(\mathbf{x}^{\prime})={\mathbf{y}}].$
-Instead of entry-level adjacency, one may consider individual-level adjacency that two inputs are adjacent if they differ at one row (i.e., an individual can decide whether to report their whole history or not).
+Instead of entry-level adjacency, one may consider individual-level adjacency that two inputs are adjacent if they differ in one row (that is, an individual can decide whether to report their entire history or not).
 Our analysis in the paper still holds after strengthening the privacy constraints from $\epsilon$ to $\epsilon/m$; please refer to Section 6 for a more detailed discussion.
-We believe the study of influence sample DP is more natural and practical than conventional node and edge DP for seeding algorithms (Kasiviswanathan et al.
+We believe that the study of influence sample DP is more natural and practical than conventional node and edge DP for seeding algorithms (Kasiviswanathan et al.
 2013, Raskhodnikova and Smith 2016).
-First, the notion of node DP allows an agent to misreport her connection in the social networks.
-However, when collecting data on network cascades, a single agent can hardly modify her connection to change the dynamics of social contagion, but rather she can decide whether to misreport her previous contagion data.
-Therefore, we argue that defining adjacent datasets based on influence samples as the input, rather than the raw social network data, is more natural and realistic.
-Moreover, we show in LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg that node-DP and edge-DP are too strict for influence maximization and hardly admit any utility guarantees.
-In Section 4 we will elaborate on the nature of our privacy protections in social network data collection and intervention design contexts.
-Give two edge-weighted undirected graphs ${\mathcal{G}}=(V,p)$ and ${\mathcal{G}}^{\prime}=(V^{\prime},p^{\prime})$, we say ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ are node adjacent if $V=V^{\prime}$ and there exists $v\in V$, $p_{e}=p_{e}^{\prime}$ for all $v\notin e$.
-Similarly, ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ are edge adjacent if $V=V^{\prime}$ and there exists $u,v\in V$ $p_{e}=p_{e}^{\prime}$ for all $e\neq(u,v)$.
+First, the notion of node DP allows an agent to misreport her social network connections.
+However, when collecting data on network cascades, a single agent can hardly modify her connections to change the contagion dynamics but instead can decide whether to misreport her contagion data.
+Therefore, we argue that defining adjacent datasets based on influence samples as input, rather than raw social network data, is more natural and realistic for contact tracing and public health applications.
+Moreover, we show in LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg that node DP and edge DP are too strict for influence maximization and hardly admit any utility guarantees.
+In Section 4, we will elaborate on the nature of our privacy protections in social network contexts, including for data collection and intervention design.
+Given two edge-weighted undirected graphs ${\mathcal{G}}=(V,p)$ and ${\mathcal{G}}^{\prime}=(V^{\prime},p^{\prime})$, we say ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ are node adjacent if $V=V^{\prime}$ and there exist $v\in V$, such that $p_{e}=p_{e}^{\prime}$ for all $e\notniFromTxfonts v$.
+Similarly, ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ are edge adjacent if $V=V^{\prime}$ and there exists $u,v\in V$ such that $p_{e}=p_{e}^{\prime}$ for all $e\neq(u,v)$.
 For $k<n$, we say that a seeding algorithm $M^{k}$ is $\epsilon$-node (or edge) DP if for any $k$, node (or edge) adjacent graphs ${\mathcal{G}},{\mathcal{G}}^{\prime}$, and output ${\mathbf{y}}$, $\Pr[M^{k}({\mathcal{G}})={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}({\mathcal{G}}^{\prime})={\mathbf{y}}]$.
-The following proposition (proved in Section 7.1) shows no node DP seeding algorithm can output a high quality seed.
-Proposition 2.5 (label = prop:node:dp:neg, name=Hardness of Influence Maximization with Node-DP)
-Given $n\geq 1$ and $\epsilon>0$, for any $\epsilon$-node-DP seeding algorithm $M^{k}({\mathcal{G}})$, there exists ${\mathcal{G}}$ of $n$ nodes so that for $k=1$ we have: $\max_{S:|S|=1}I_{\mathcal{G}}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}({\mathcal{G}})}[I_{\mathcal{G}}({\mathbf{y}})]\geq\frac{n}{4}-(e^{\epsilon}+1).$
+The following proposition (proved in Section 7.1) shows that seeding algorithms with node DP guarantees cannot produce high-quality seeds.
+Proposition 2.5 (label = prop:node:dp:neg, name=Hardness of Influence Maximization with Node DP)
+Given $n\geq 1$ and $\epsilon>0$, for any $\epsilon$ node-DP seeding algorithm $M^{k}({\mathcal{G}})$, there exists ${\mathcal{G}}$ of $n$ nodes so that for $k=1$ we have: $\max_{S:|S|=1}I_{\mathcal{G}}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}({\mathcal{G}})}[I_{\mathcal{G}}({\mathbf{y}})]\geq\frac{n}{4}-(e^{\epsilon}+1).$
 We use the fact that a star and a totally isolated graph are different by one node’s connection.
-A node-DP algorithm needs to output similar seed sets on these two graphs, and thus returns low quality seeds.
-Based on similar intuition, the following proposition (proved in Section 7.2) shows no edge-DP seeding algorithm can output a high quality seed either.
+A node-DP algorithm needs to output similar seed sets on these two graphs and thus returns low-quality seeds.
+Based on a similar intuition, the following proposition (proved in Section 7.2) shows that no edge-DP seeding algorithm can output a high-quality seed either.
 Both results can be strengthened to exclude the possibility of useful algorithms with “approximate” $(\epsilon,\delta)$-DP guarantees as well.
 The latter is a relaxation of the defining constraint in Equation 1 to allow for an additive term $\delta$; see Section 7.2.
-Proposition 2.6 (label = prop:edge:dp:neg, name=Hardness of Influence Maximization with Edge-DP)
-For all $n\geq 1$ and $\epsilon>0$, there exists $\alpha_{\epsilon}\geq\frac{1}{1000e^{\epsilon}}$ so that any $\epsilon$-edge-DP seeding algorithm $M^{k}({\mathcal{G}})$, there exists graph ${\mathcal{G}}$ on $n$ nodes so that for $k=1$ we have: $(1-1/e)\textstyle\max_{S:|S|=1}I_{\mathcal{G}}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}({\mathcal{G}})}[I_{\mathcal{G}}({\mathbf{y}})]\geq\alpha_{\epsilon}n.$
-Compared to LABEL:thm:original, the above results show that no node-DP (or edge-DP) algorithm can perform close to $(1-1/e)\mbox{\sf OPT}$ even with the full knowledge of ${\mathcal{G}}$.
-In Section 7.2 (Proposition 7.1), we show that similar lower bounds hold even if we consider an additive relaxation of the node- and edge-DP requirements, allowing the $\Pr[M^{k}({\mathcal{G}})={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}({\mathcal{G}}^{\prime})={\mathbf{y}}]$ condition to be violated up to a $\delta$ additive factor.
+Proposition 2.6 (label = prop:edge:dp:neg, name=Hardness of Influence Maximization with Edge DP)
+For all $n\geq 1$, $\epsilon>0$, and any $\epsilon$ edge-DP seeding algorithm $M^{k}({\mathcal{G}})$, there exists a graph ${\mathcal{G}}$ on $n$ nodes so that for $k=1$ and some $\alpha_{\epsilon}\geq\frac{1}{1000e^{\epsilon}}$ we have: $(1-1/e)\textstyle\max_{S:|S|=1}I_{\mathcal{G}}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}({\mathcal{G}})}[I_{\mathcal{G}}({\mathbf{y}})]\geq\alpha_{\epsilon}n.$
+Compared to LABEL:thm:original, the above results show that node-DP (or edge-DP) algorithms cannot perform close to $(1-1/e)\mbox{\sf OPT}$ even with full knowledge of ${\mathcal{G}}$.
+In Section 7.2 (Proposition 7.1), we show that similar lower bounds hold even if we consider additive relaxations of the node and edge DP, allowing the $\Pr[M^{k}({\mathcal{G}})={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}({\mathcal{G}}^{\prime})={\mathbf{y}}]$ condition to be violated up to a $\delta$ additive factor.
 In this paper, we will design influence sample DP algorithms that perform comparable to Algorithm 1.
-Finally, note that node-DP implies edge-DP, but influence sample DP is not comparable to node- or edge-DP because the algorithms take different inputs.
+Finally, note that node DP implies edge-D , but influence sample DP is not comparable to node or edge DP because the algorithms take different inputs.
 Two primitive DP algorithms.
-Our algorithms will make use of the following two general DP mechanisms (we use algorithm and mechanism interchangeably): exponential mechanism (McSherry and Talwar 2007, Dwork et al.
+Our algorithms make use of the following two general DP mechanisms (we use algorithm and mechanism interchangeably): exponential mechanism (McSherry and Talwar 2007, Dwork et al.
 2014), and randomized respond mechanism (Warner 1965, Evfimievski et al.
 2003).
-First, to achieve $\epsilon$-DP, given a finite set (range), $\mathcal{R}$, the exponential mechanism is defined with respect to a scoring function $u:\{0,1\}^{n\times m}\times\mathcal{R}\to{\mathbb{R}}$ and its sensitivity $\Delta(u):=\max_{r\in\mathcal{R},{\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}}|u({\mathbf{x}},r)-u({\mathbf{x}}^{\prime},r)|$.
-The exponential mechanism $\mathcal{M}^{\epsilon}_{E}({\mathbf{x}},u,\mathcal{R})$ outputs an element $r\in\mathcal{R}$ with probability proportional to $\exp\left(\frac{\epsilon u({\mathbf{x}},r)}{2\Delta(u)}\right)$ as defined in Dwork et al.
+First, to achieve $\epsilon$-DP, given a finite set (range), $\mathcal{R}$, exponential mechanism is defined with respect to a scoring function $u:\{0,1\}^{n\times m}\times\mathcal{R}\to{\mathbb{R}}$ and its sensitivity $\Delta(u):=\max_{r\in\mathcal{R},{\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}}|u({\mathbf{x}},r)-u({\mathbf{x}}^{\prime},r)|$.
+The exponential mechanism $\mathcal{M}^{\epsilon}_{E}({\mathbf{x}},u,\mathcal{R})$ produces an element $r\in\mathcal{R}$ with a probability proportional to $\exp\left(\frac{\epsilon u({\mathbf{x}},r)}{2\Delta(u)}\right)$ as defined in Dwork et al.
 (2014).
-Moreover, as shown in the following theorem, for a fixed dataset ${\mathbf{x}}$ exponential mechanism’s output almost maximizes the scoring function $u({\mathbf{x}},r)$ and can be interpreted as a DP soft max.
+Moreover, as shown in the following theorem, for a fixed data set ${\mathbf{x}}$, the output of the exponential mechanism almost maximizes the scoring function $u({\mathbf{x}},r)$ and can be interpreted as a softmax that satisfies the DP randomization requirement.
 Theorem 2.7 (label = thm:exp:utility, name = Theorem 3.11 of Dwork et al. (2014) — Exponential Mechanism)
 Consider a data set ${\mathbf{x}}\in\{0,1\}^{n\times m}$, privacy budget $\epsilon>0$, a finite set $\mathcal{R}$, and a scoring function $u:\{0,1\}^{n\times m}\times\mathcal{R}\to{\mathbb{R}}$ with sensitivity $\Delta(u)$.
-The exponential mechanism $\mathcal{M}^{\epsilon}_{E}({\mathbf{x}},u,\mathcal{R})$ that outputs an element $r\in\mathcal{R}$ with probability proportional to $\exp\left(\frac{\epsilon u({\mathbf{x}},r)}{2\Delta(u)}\right)$ is $\epsilon$-ISDP and for all $t>0$, we have:
+The exponential mechanism $\mathcal{M}^{\epsilon}_{E}({\mathbf{x}},u,\mathcal{R})$ that produces an element $r\in\mathcal{R}$ with a probability proportional to $\exp\left(\frac{\epsilon u({\mathbf{x}},r)}{2\Delta(u)}\right)$ is $\epsilon$-ISDP and for all $t>0$, we have:
 $\Pr\left[u({\mathbf{x}},\mathcal{M}^{\epsilon}_{E}({\mathbf{x}},u,\mathcal{R}))\leq\max_{r\in\mathcal{R}}u({\mathbf{x}},r)-\frac{2\Delta(u)}{\epsilon}(\ln|\mathcal{R}|+t)\right]\leq e^{-t},$
@@
 The above exponential mechanism ensures that the output does not reveal too much information about the privacy-critical input data, limiting how much an adversary can infer about the input data according to Definition 2.4.
-However, depending on data curation requirements (who can access the exact user data), protection against output-based inferences may not be sufficient and we may need to limit the algorithm’s access to exact input data.
-Such privacy guarantees can be formulated as local DP where each individual’s data is privatized before running the algorithms.
-One standard locally DP algorithm is the randomized response mechanism (Evfimievski et al.
-2003).
-Specifically, an $\epsilon$-DP randomized response mechanism guarantees privacy by independently flipping each bit of the input dataset (influence sample collection) with probability $1/(e^{\epsilon}+1)$.
+However, depending on data curation requirements (who can access the exact user data), protection against output-based inferences may not be sufficient, and we may need to limit the algorithm’s access to exact input data.
+Such privacy guarantees can be formulated as local DP where the data of each individual are privatized before running the algorithms.
+A standard locally DP algorithm is randomized response mechanism (Evfimievski et al.
+2003): an $\epsilon$-DP randomized response mechanism guarantees privacy by independently flipping each bit of the input data set (influence sample collection) with probability $1/(e^{\epsilon}+1)$.
 3 Differentially Private Seeding Algorithms
 In this section, we provide two DP seeding algorithms that satisfy the DP requirements either centrally or locally.
-First in Subsection 3.1, we use an exponential mechanism that ensures the output of the seed selection is centrally DP.
-Next in Subsection 3.2, we use the randomized response algorithm to achieve local DP according to the criteria set forth in Definition 2.4.
+First in Section 3.1, we use an exponential mechanism that ensures that the seed selection output is centrally DP.
+Next in Section 3.2, we use the randomized response algorithm to achieve local DP according to the criteria set forth in Definition 2.4.
 LABEL:thm:exp:acc and LABEL:thm:rr:acc provide our sample complexity bounds in each case to achieve $\epsilon$-ISDP with a $(1-1/e)\mbox{\sf OPT}-\alpha n$ performance guarantee.
 In particular, we observe that the sample complexity bound for achieving local DP is much harsher at $O(k^{3}\epsilon^{-2k^{2}}\ln n/\alpha^{2})$ compared to $O(k\ln n/(\alpha\epsilon))$ for central DP.
-This exponential separation of sample complexity dependence on $k$ is reflected in our simulation results in Section 5, where exponential mechanism shows a faster improvement in performance with increasing number of influence samples and its output has a significantly higher expected spread size when the privacy budget ($\epsilon$) is low.
+This exponential separation of the dependency of sample complexity on $k$ is reflected in our simulation results in Section 5, where the exponential mechanism shows a faster improvement in performance with an increasing number of influence samples and its output has a significantly higher expected spread size when the privacy budget ($\epsilon$) is low.
 3.1 Central DP with Exponential Mechanism
-In this section, we design a DP seeding algorithm by using exponential mechanism and modifying Algorithm 1.
+In this section, we design a DP seeding algorithm using the exponential mechanism and modifying Algorithm 1.
 To protect privacy, instead of Algorithm 2, we use an exponential mechanism that differentially outputs a softmax (LABEL:thm:exp:utility).
-Formally, given $\epsilon>0$ and integer $k$ we define a seeding algorithm $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ that runs Algorithm 1 but replaces the single seed selection steps of Algorithm 2 by its exponential mechanism variant, Algorithm 3 $\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}},S)$.
-Note that Algorithm 3 is an exponential mechanism using $I_{\mathbf{x}}(\cdot|S)$ as the scoring function.
+Formally, given $\epsilon>0$ and the integer $k$ we define a seeding algorithm $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ that runs Algorithm 1 but replaces the single seed selection steps of Algorithm 2 by its variant of the exponential mechanism Algorithm 3 $\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}},S)$.
+Note that Algorithm 3 is an exponential mechanism that uses $I_{\mathbf{x}}(\cdot|S)$ as the scoring function.
 Algorithm 3 $\epsilon$-ISDP single seed selection, $\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}},S)$
@@
 Note that a naive algorithm that runs the exponential mechanism directly on all $k$ seeds at once is not efficient because the size of the output space is combinatorially large $\Omega(n^{k})$.
-To ensure our algorithm is efficient, we sequentially run the exponential mechanism $k$ times to estimate and choose seeds similar to the greedy algorithm of Eckles et al.
-(2022), reducing their time complexity to $O(nmk)$.
-LABEL:thm:exp:inf shows that our exponential mechanism for single seed selection, $\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}},S)$, is $\epsilon$-ISDP (Definition 2.4), and provides a tail bound on the approximation error of maximizing $I_{\mathbf{x}}$.
-Its proof in Section 7.3 follows from the facts that $I_{\mathbf{x}}$ is a submodular function, and the exponential mechanism in Algorithm 3 can be seen as a noisy maximum function which implements an approximate greedy step by LABEL:thm:exp:utility.
+To ensure that our algorithm is efficient, we sequentially run the exponential mechanism $k$ times to estimate and choose seeds similar to the greedy algorithm of Eckles et al.
+(2022), but reducing their time complexity to $O(nmk)$.
+LABEL:thm:exp:inf shows that our modified greedy algorithm ($\mathcal{M}_{exp}^{k,\epsilon}({\mathbf{x}})$) that applies the exponential mechanism ($\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}},S)$) for single seed selection $k$ times, is $\epsilon$-ISDP following Definition 2.4, and also provides a tail bound on the approximation error for maximization of $I_{\mathbf{x}}$ using $\mathcal{M}_{exp}^{k,\epsilon}({\mathbf{x}})$.
+Its proof in Section 7.3 follows from the facts that $I_{\mathbf{x}}$ is a submodular function, and the exponential mechanism in Algorithm 3 can be seen as a noisy maximum function that implements an approximate greedy step in LABEL:thm:exp:utility.
 Theorem 3.1 (label = thm:exp:inf)
@@
 Theorem 3.2 (label = thm:exp:acc)
-Given graph ${\mathcal{G}}$ on $n\geq 2$ nodes and $k\leq n$, for any $0<\alpha\leq 1$ and $\epsilon>0$, $\mathcal{M}^{k,\epsilon}_{exp}(\cdot)$ that runs Algorithm 1 but replaces the single seed selection steps of Algorithm 2 by $\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}},S)$, its exponential mechanism variant in Algorithm 3, is $\epsilon$-ISDP and its output infects at least $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes with high probability, using at least $m=\max\{\frac{12}{\alpha\epsilon},\frac{9}{\alpha^{2}}\}k\ln n$ influence samples in $O(knm)$ run time.
-Note that if we are not concerned with privacy and set $\epsilon\gg\alpha$, then the number of samples to achieve $\alpha n$ additive error is of order $\alpha^{-2}k\ln n$, which improves upon the $k^{2}(\ln n/\alpha)/\alpha^{3}$ bound of Eckles et al.
+Given graph ${\mathcal{G}}$ on $n\geq 2$ nodes and $k\leq n$, for any $0<\alpha\leq 1$ and $\epsilon>0$, $\mathcal{M}^{k,\epsilon}_{exp}(\cdot)$ that runs Algorithm 1 but replaces the single seed selection steps of Algorithm 2 by its exponential mechanism variant in Algorithm 3, $\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}},S)$, is $\epsilon$-ISDP and its output infects at least $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes with high probability, using at least $m=\max\{\frac{12}{\alpha\epsilon},\frac{9}{\alpha^{2}}\}k\ln n$ influence samples in $O(knm)$ run time.
+Note that if we are not concerned with privacy and set $\epsilon\gg\alpha$, then the number of samples to achieve the additive error $\alpha n$ is of the order $\alpha^{-2}k\ln n$, which improves the $k^{2}(\ln n/\alpha)/\alpha^{3}$ bound of Eckles et al.
 (2022).
 Moreover, our result does not violate Theorem 3 of Eckles et al.
-(2022) that shows $\Omega(k^{2})$ influence samples are required to achieve $(1-1/e)\mbox{\sf OPT}-\alpha n$ performance bound for all $\alpha<(1-1/e)/k$, because our algorithm also needs $\Omega(k^{2})$ samples when $\alpha<1/k$.
+(2022) which shows that $\Omega(k^{2})$ influence samples are required to achieve the $(1-1/e)\mbox{\sf OPT}-\alpha n$ performance guarantee for all $\alpha<(1-1/e)/k$, because our algorithm also needs $\Omega(k^{2})$ samples when $\alpha<1/k$.
 3.2 Local DP with Randomized Response
 We can achieve a stronger notion of privacy, local DP, by adding noise to the input data before running our (or any other) algorithms.
-Line 3 of Algorithm 4 adds the requisite amount of noise to achieve local $\epsilon$-ISDP in the input dataset ${\mathbf{x}}$ — flipping each entry of ${\mathbf{x}}$ independently with probability $\rho_{\epsilon}=1/(1+e^{\epsilon})$ to generate the noisy input $\tilde{{\mathbf{x}}}$.
+Line 3 of Algorithm 4 adds the necessary amount of noise to achieve local $\epsilon$-ISDP in the input dataset ${\mathbf{x}}$ — flipping each entry of ${\mathbf{x}}$ independently with probability $\rho_{\epsilon}=1/(1+e^{\epsilon})$ to generate noisy input $\tilde{{\mathbf{x}}}$.
 However, the added input noise will bias our estimate of each candidate seed’s quality; in particular, $\operatorname*{\mathbb{E}}_{\tilde{{\mathbf{x}}}}[I_{\tilde{{\mathbf{x}}}}(S|T)]$ is not necessarily equal to $I_{\mathcal{G}}(S|T)$ for a pair of subsets $S,T\subset V$ and we no longer have the benefit of applying Proposition 2.1 directly.
 Rather in Algorithm 5 we propose a post-processing of the noisy input data $\tilde{{\mathbf{x}}}$ to derive an unbiased estimate of a seed set’s quality.
-Importance of proper post-processing steps in the analysis of DP data are highlighted in other applications, e.g., census DP data release, where ignoring them can lead to biased decision-making and sub-optimal outcomes (Abowd et al.
+The importance of proper post-processing steps in the analysis of DP data is highlighted in other applications, e.g., in the release of census DP data, where ignoring them can lead to biased decision-making and suboptimal results (Abowd et al.
 2019, Ashmead et al.
 2019, Zhu et al.
-2022, 2021).
+2022, 2021, 2023).
 The added variability of the estimators due to the DP noise can also degrade our estimation performance (even if not biased), which is a second concern in local DP algorithm design and opens a venue to explore alternative post-processing methods with potentially superior bias-variance tradeoff.
-In the remainder of this section, we first introduce Algorithms 4 and 5 and then offer LABEL:thm:rr:acc for their privacy and utility guarantees, addressing both concerns about the bias and variance of decision making with local DP influence samples.
+In the remainder of this section, we first introduce Algorithms 4 and 5 and then offer LABEL:thm:rr:acc for their privacy and utility guarantees, addressing both concerns about bias and variance of decision making with local DP influence samples.
 Algorithm 4 Local $\epsilon$-ISDP seeding algorithm, $\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}})$
@@
 By Proposition 2.1, the influence function $I_{\mathcal{G}}(S)$ can be estimated from the reachable set, $n\Pr[S\cap x\neq\emptyset]$.
-But to achieve local DP, we can only observe the perturbed, reachable set $S\cap\tilde{x}$.
+However, to achieve local DP, we can only observe the perturbed, reachable set $S\cap\tilde{x}$.
 In other words, let $\tilde{z}_{S}:=|S\cap\tilde{x}|$ and ${z}_{S}:=|S\cap{x}|$ be two random variables.
-If we can estimate $\Pr[z_{S}=0]$ from our observed samples of $\tilde{z}_{S}$, then we can estimate the size of the reachable set from $I_{\mathcal{G}}(S)=n\Pr[z_{S}>0]$ — using Proposition 2.1 — and then run the standard greedy algorithm for seed selection based on the estimated values.
+If we can estimate $\Pr[z_{S}=0]$ from our observed samples of $\tilde{z}_{S}$, then we can estimate the size of the reachable set from $I_{\mathcal{G}}(S)=n\Pr[z_{S}>0]$ — using Proposition 2.1 — and then run the standard greedy algorithm for seed selection based on estimated values.
 Now, we discuss how to estimate $\Pr[z_{S}=0]$ from samples of $\tilde{z}_{S}$.
@@
 Note that $S\cap x$ is the set of elements in $S$ that is also in the influence sample $x$.
-Hence, $S\cap\tilde{x}$ can be partitioned into (1) the subset $(S\cap\tilde{x})\cap x$ of elements of $x$ that remain in $S$ after perturbation, and (2) the subset $(S\cap\tilde{x})\setminus x$ of the elements of $\tilde{x}$ that join $S$ after the perturbation.
+Hence, $S\cap\tilde{x}$ can be partitioned into (1) the subset $(S\cap\tilde{x})\cap x$ of elements of $x$ that remain in $S$ after perturbation and (2) the subset $(S\cap\tilde{x})\setminus x$ of the elements of $\tilde{x}$ that join $S$ after perturbation.
 The size of the first subset is a binomial random variable $Bin(|S\cap x|,1-\rho)$, and the size of the second subset is $Bin(|S\setminus x|,\rho)$.
 Therefore, given $|S|=l$ for all $a$ and $b$, the probability of $\tilde{z}_{S}=b$ conditional on $z_{S}=a$ is $\Pr[\tilde{z}_{S}=a|z_{S}=b]=\Pr[Bin(b,1-\rho)+Bin(l-b,\rho)=a],$ for all $a,b=0,\dots,l$ where $Bin(b,1-\rho)$ and $Bin(l-b,\rho)$ are independent.
-Additionally, if we set matrix $C=C_{\rho,l}\in{\mathbb{R}}^{(l+1)\times(l+1)}$ be a likelihood function where $C(a,b)=C_{\rho,l}(a,b)=\Pr[\tilde{z}_{S}=a|z_{S}=b]$, then we have:
+Furthermore, if we set the matrix $C=C_{\rho,l}\in{\mathbb{R}}^{(l+1)\times(l+1)}$ as the likelihood function where $C(a,b)=C_{\rho,l}(a,b)=\Pr[\tilde{z}_{S}=a|z_{S}=b]$, then we have the following:
 $\begin{pmatrix}\Pr[\tilde{z}_{S}=0]\\ \Pr[\tilde{z}_{S}=1]\\ \vdots\\ \Pr[\tilde{z}_{S}=l]\end{pmatrix}=C_{\rho,l}\begin{pmatrix}\Pr[{z}_{S}=0]\\ \Pr[{z}_{S}=1]\\ \vdots\\ \Pr[{z}_{S}=l]\end{pmatrix}.$ (2)
-We can write out each entry explicitly as follows:
+We can write each entry of the $C_{\rho,l}$ matrix explicitly as follows:
 $C_{\rho,l}(a,b)=\sum_{\ell=\max(0,b-a)}^{\min(l-a,b)}\binom{b}{\ell}\binom{l-b}{a-b+\ell}\rho^{a-b+2\ell}(1-\rho)^{l-a+b-2\ell}.$ (3)
-To see why, condition on $\ell$ of the $z_{S}=b$ nodes in $S\cap x$ being flipped, leaving $b-\ell$ nodes in $S\cap x\cap\tilde{x}$.
-To get $\tilde{z}_{S}=a$ nodes in $S\cap\tilde{x}$, the remaining $a-b+\ell$ nodes should come from perturbing the status of the $l-b$ nodes belonging to $S\setminus x$.
-Note that our indexes $a$ and $b$ start at 00 instead of $1$.
-In the rest of the section, we will adopt this convention.
-If matrix $C_{\rho,l}$ is invertible, then we can have an unbiased estimation of $\Pr[z_{S}>0]$ using Algorithm 5, below.
-LABEL:prop:feasible, proved in Section 7.5, shows that $C_{\rho.l}$ is invertible for $\rho>1/2$ and thus completes the specification of our randomized response algorithm.
-Algorithm 5 Unbiased estimate for number of reachable set $J_{m}(S)$
+To see why, consider the event that $\ell$ of the $z_{S}=b$ nodes in $S\cap x$ are flipped, leaving the $b-\ell$ nodes in $S\cap x\cap\tilde{x}$.
+To obtain $\tilde{z}_{S}=a$ nodes in $S\cap\tilde{x}$, the remaining $a-b+\ell$ nodes should come from perturbing the status of the $l-b$ nodes belonging to $S\setminus x$.
+Note that our indices ($a$ and $b$) start at 00 instead of $1$.
+In the remainder of the section, we will adopt this convention.
+If the matrix $C_{\rho,l}$ is invertible, then we can have an unbiased estimate of $\Pr[z_{S}>0]$ using Algorithm 5, below.
+LABEL:prop:feasible, proved in Section 7.5, shows that $C_{\rho,l}$ is invertible for $\rho>1/2$ and thus completes the specification of our randomized response algorithm.
+Algorithm 5 Unbiased estimate for the size of the reachable set $J_{m}(S)$
 1:$m$ perturbed influence samples $\tilde{{\mathbf{x}}}=(\tilde{x}^{1},\ldots,\tilde{x}^{m})$, their perturbation probability $\rho$, and a candidate set $S\subset V$ with size $|S|=l$.
@@
 Proposition 3.3 (label = prop:feasible, name = Feasibility of Unbiased Post-processing)
-Given any positive integer $l\geq 1$ and $\rho\in[0,1/2)$, matrix $C_{\rho,l}$ is invertible.
-Note that when $\rho=1/2$, $\tilde{{\mathbf{x}}}$ consists of uniform random bits, and we cannot recover $\Pr[z_{S}>0]$ from the perturbed influence samples, and LABEL:prop:feasible shows that we can recover $\Pr[z_{S}>0]$ for any $\rho<1/2$.
+Given any positive integer $l\geq 1$ and $\rho\in[0,1/2)$, the matrix $C_{\rho,l}$ is invertible.
+Note that when $\rho=1/2$, $\tilde{{\mathbf{x}}}$ consists of uniform random bits and we cannot recover $\Pr[z_{S}>0]$ from perturbed influence samples.
+LABEL:prop:feasible shows that we can recover $\Pr[z_{S}>0]$ for any $\rho<1/2$.
 Theorem 3.4 (label = thm:rr)
@@
 The above theorem, proved in Section 7.6, provides a tail bound on the approximation error of $\mathcal{M}_{loc}^{k,\epsilon}$ given a fixed number of influence samples.
-We can convert the above result to a high probability error bound by choosing $m=\Omega({k^{3}V_{\rho_{\epsilon},k}^{2}\ln n}/{\alpha^{2}})$ which for $k$ constant and $\epsilon\gg\alpha$ improves on the $1/\alpha^{3}$ sample complexity in LABEL:thm:original (complete proof in Section 7.7):
+We can convert the above result to a high probability error bound by choosing $m=\Omega({k^{3}V_{\rho_{\epsilon},k}^{2}\ln n}/{\alpha^{2}})$ which for $k$ constant and $\epsilon\gg\alpha$ improves the sample complexity ($1/\alpha^{3}$) in LABEL:thm:original (complete proof in Section 7.7):
 Theorem 3.5 (label = thm:rr:acc)
 Given graph ${\mathcal{G}}$ with $n\geq 2$ nodes and $k\leq n$, for any $0<\alpha\leq 1$ and $\epsilon>0$, $\mathcal{M}_{loc}^{k,\epsilon}$ is locally $\epsilon$-ISDP and its output infects at least $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes with high probability, using at least $m=O({k^{3}V_{\rho_{\epsilon},k}^{2}\ln n}/{\alpha^{2}})$ influence samples in $O(nk^{4}+kn^{2}m)$ run time, where $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})$.
-Note that $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})$ in LABEL:thm:rr and LABEL:thm:rr:acc can be simplified as follows: $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})=O((1/2-1/(e^{\epsilon}+1))^{-k^{2}})=O([2(e^{\epsilon}+1)/(e^{\epsilon}-1)]^{k^{2}})$.
-Therefore $V_{\rho_{\epsilon},k}=O(\epsilon^{-k^{2}})$ as $\epsilon\to 0$ and $V_{\rho_{\epsilon},k}=O(2^{k^{2}})$ as $\epsilon\to\infty$.
+We conclude this section by considering two scenarios: the non-private regime $\epsilon\to\infty$ and the full-privacy regime $\epsilon\to 0$.
+First, in the non-private regime ($\epsilon\to\infty$), $\rho_{\epsilon}\to 0$, which causes the matrix $C_{\rho,l}$ and its inverse to converge to the identity matrix, so we can take $V_{\rho_{\epsilon},k}$ to be a constant arbitrarily close to one.
+As a result, $m=O\left(\frac{k^{3}\ln n}{\alpha^{2}}\right)$ samples are sufficient, which is comparable to LABEL:thm:original with an additional factor of $k$, but improves the dependence on $\alpha$.
+We may reduce the dependence on $k$ from $k^{3}$ to a smaller order, using techniques in adaptive data analysis Bassily et al.
+(2016).
+Informally, our analysis requires that the query error in any possible seed set $S$ with $|S|\leq k$ is small enough (Lemma 7.5) so that the output of the greedy algorithm is good enough.
+However, we may not need such stringent conditions, because the greedy algorithm does not necessarily query all possible seed sets.
+Alternatively, adaptive data analysis provides a procedure to support adaptive queries from the greedy algorithm with a better error bound.
+[^1] Interestingly, the procedure in Bassily et al.
+(2016) adds Gaussian noise to each query and achieves a small error guarantee using techniques from an exponential mechanism similar to  LABEL:thm:exp:acc.
+Furthermore, in the non-private regime ($\epsilon\to\infty$), our exponential mechanism and the local DP seeding algorithm are identical, allowing us to use the analysis in LABEL:thm:exp:inf and derive an $O\left(\frac{k\ln n}{\alpha^{2}}\right)$ upper bound for the required number of influence samples.
+On the other hand, if $\epsilon\to 0$, $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})$ in LABEL:thm:rr and LABEL:thm:rr:acc can be simplified as follows: $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})=O((1/2-1/(e^{\epsilon}+1))^{-k^{2}})=O([2(e^{\epsilon}+1)/(e^{\epsilon}-1)]^{k^{2}})=O(\epsilon^{-k^{2}})$.
 4 The Nature of Our Privacy Protections in Social Networks
-From the perspective of privacy guarantees, exponential mechanism (central DP) only protects the input data against the algorithmic revelations of the output (seeded nodes), giving the algorithm itself unhindered access to user data.
-On the other hand, the randomized response mechanism (local DP) provides a stronger privacy notion that protects the input data directly against any (adversarial) misuse, regardless of – and including — their use for seeding.
-Our results address both central and local notions of privacy, providing performance (utility) and sample complexity guarantees for influence maximization algorithms whose outputs/inputs are randomized at a desired level to give central/local DP protection given our input data structure, i.e., influence samples.
-In our framework, we explicitly establish a budget $(\epsilon)$ for privacy leaks through the observed network cascades.
-This budget is a quantifiable limit on how much privacy/security risk one is willing to tolerate across all queries, not just a single query of the data.
-The latter highlights yet another way, in which our formulation of privacy budget is distinct from the setting of node- and edge-DP for graph input data structures.
-While algorithmically tractable, practice-wise our notion of adjacency protects each data donor against revelations of their contributions to any given cascade.
-While a donor may be willing to donate their contagion history, there may be some cascades in that history that they feel uncomfortable about (revealing sensitive or compromising information).
-Or the donors may dynamically contribute their contagion data where previous donations are irrevocable.
-Our notion of influence sample differential privacy provides donors with plausible deniability at any fixed cascade.
-This way donors are incentivized to contribute their entire set of $m$ cascades because we can bound their privacy risks from information leakage at any specific cascade in the set of size $m$, either centrally or locally - depending on the data curation environment - using our Algorithms 3 and 4.
-As long as the number of sensitive cascades is much smaller than $m$, we believe this notion of adjacency provide meaningful privacy protections.
-This is reflected in the requisite randomization in our greedy exponential mechanism (Algorithm 3): to achieve $\epsilon-$DP, each node $v$ is chosen with probability proportional to $\exp\left(\frac{\epsilon m}{2n}I_{\mathbf{x}}(v|S)\right)$, where $I_{\mathbf{x}}(v|S)$ is the marginal benefit of adding $v$ to the current seed set $S$ and it is estimated from the input influence samples ${\mathbf{x}}$.
+From the perspective of privacy guarantees, the exponential mechanism (central DP) only protects the input data against the algorithmic revelations of the output (seeded nodes), giving the algorithm itself unhindered access to user data.
+On the other hand, the randomized response mechanism (local DP) provides a stronger privacy notion that protects the input data directly against any (adversarial) misuse, regardless of – and including – their use for seeding.
+Our results address both central and local notions of privacy, providing performance (utility) and sample complexity guarantees for influence maximization algorithms whose outputs/inputs are randomized at a desired level to give central/local DP protection given the input data structure of influence samples.
+In our framework, we explicitly establish a privacy budget $(\epsilon)$ to limit privacy leaks through observed influence cascades, such as those collected during contact tracing for HIV prevention.
+This budget quantifies the amount of privacy or security risk that can be tolerated across all queries, rather than focusing on a single query of the data.
+This approach distinguishes our privacy budget formulation from node and edge differential privacy (DP) commonly applied to graph data structures.
+From a privacy standpoint, our notion of adjacency protects each data donor by ensuring that their contribution to any given cascade remains private.
+Although a data donor may be willing to share their history of sexual activity as part of an ongoing contact tracing effort, there may be certain cascades in that history that they prefer to keep private.
+In addition, donors can dynamically contribute their contagion data knowing that prior contributions cannot be revoked.
+Our definition of influence sample differential privacy offers donors plausible deniability for any individual cascade.
+This ensures that they are incentivized to provide their entire set of $m$ cascades, knowing that their privacy risk in each cascade is bounded, centrally or locally, depending on the data curation environment.
+These privacy guarantees are implemented in our Algorithms 3 and 4, ensuring that even in sensitive settings such as HIV prevention, participants’ data is handled securely while allowing effective interventions.
+As long as the number of sensitive cascades is much smaller than $m$, this notion of adjacency provides meaningful central privacy protections.
+This is reflected in the randomization required in our greedy exponential mechanism (Algorithm 3): To achieve $\epsilon-$DP, each node $v$ is chosen with a probability proportional to $\exp\left(\frac{\epsilon m}{2n}I_{\mathbf{x}}(v|S)\right)$, where $I_{\mathbf{x}}(v|S)$ is the marginal benefit of adding $v$ to the current seed set $S$ and is estimated from the input influence samples ${\mathbf{x}}$.
 As $m\to 0$, this “softmax” converges to a uniform random choice, completely ignoring the input influence samples ${\mathbf{x}}$.
 On the other hand, the randomized response mechanism achieves $\epsilon-$DP locally by flipping the status of nodes in each cascade independently at random with probability $1/(1+e^{\epsilon})$.
-While the randomized response noise level is determined independently of the sample size $m$, the latter scales up as $m=O({k^{3}\epsilon^{-2k^{2}}\ln n}/{\alpha^{2}})$ with the decreasing privacy budget $(\epsilon)$ to sustain algorithmic performance in the privacy-critical regime ($\epsilon\to 0$).
+Although the level of randomized response noise is determined independently of the sample size $m$, the latter scales as $m=O({k^{3}\epsilon^{-2k^{2}}\ln n}/{\alpha^{2}})$ with the decreasing privacy budget $(\epsilon)$ to maintain algorithmic performance in the privacy-critical regime ($\epsilon\to 0$).
 As a concrete case, consider the choice of community leaders to undergo training for a public health or education intervention.
-To identify and target community leaders for treatment, the interventionist can survey the social network connections among the population.
+To identify and target community leaders for treatment, the interventionist can survey social network connections in the population.
 For example, Banerjee et al.
-(2013) survey individuals to collect social network data along 12 dimensions, including whom they get medical advice from or give medical advice to, relative and non-relatives with whom they socialize, and people with whom they exchange goods.
-While the community leaders may consent to and be properly compensated for their time in treatment, their social contacts face a privacy risk in having one of their associates chosen.
-As such, people may refrain from contributing to social network data collection or refuse to answer survey questions, especially when association with the treated individuals can pose harm, such as stigma or retaliation.
-For example, to prevent school bullying a few students may be recruited into treatment programs — e.g., by encouraging them to take a public stance against conflict (Paluck et al.
-2016) — if the treated individuals are chosen based on their social network positions, and even if they are protected in treatment, their friends and contacts can face retaliation outside the treatment.
-Our central privacy protections in Section 3.1 allows the network participants to make a credible claim that their network associates are recruited into treatment by random choice, irrespective of their associations or contribution to the collected network cascade data.
-In other related contexts, the nature of collected data can itself be compromising, irrespectively of how individuals are selected into treatments; e.g., surveillance and control of sexually transmitted disease and HIV may entail surveys of sexual contacts or needle sharing networks among people who inject drugs (Fairley et al.
+(2013) survey individuals to collect data about their social networks in 12 dimensions, including whom they receive medical advice from or give medical advice to, relative and non-relatives with whom they socialize, and people with whom they exchange goods.
+Although community leaders may consent to treatment and be compensated for their time, their social contacts face a privacy risk in their association with treated individuals.
+As such, people may refrain from contributing to the collection of social network data or refuse to answer survey questions, especially when association with treated individuals can pose harm, such as stigma or retaliation.
+For example, to prevent school bullying, a few students can be recruited into treatment programs, e.g., by encouraging them to take a public stance against conflict (Paluck et al.
+2016).
+If treated individuals are chosen based on their social network positions and even if they are protected in treatment, their friends and contacts can face retaliation outside the treatment.
+Our central privacy protections in Section 3.1 allow network participants to make a credible claim that their network associates are recruited into treatment by random choice, regardless of their associations or contribution to the collected network cascade data.
+In other related contexts, the nature of the collected data can itself be compromising, regardless of how individuals are selected into treatments; e.g., surveillance and control of sexually transmitted diseases and HIV can involve surveys of sexual contacts or needle sharing networks among people who inject drugs (Fairley et al.
 2022, Endo et al.
 2022).
-Administering such privacy-critical data should entail local privacy protections, whereby individual data is randomized as they are collected before any intervention designs or allocation decisions.
-Our local privacy guarantees in Section 3.2 are suited for such cases where individual contagion data may be collected for intervention design, but needs to be protected locally at the collection phase.
-In a similar vein, social media platforms rely on social network data and information from past diffusions to organize their ad campaign and improve the efficacy of targeted advertising.
+Administering such privacy-critical data should entail local privacy protections, by which individual data are randomized as they are collected before any intervention designs or allocation decisions.
+Our local privacy guarantees in Section 3.2 are suitable for such cases where individual contagion data may be collected for intervention design, but must be protected locally in the collection phase.
+In a similar vein, social media platforms rely on social network data and information from previous diffusions to organize their ad campaign and improve the efficacy of targeted advertising.
 Ensuring that these network operations do not violate users’ privacy is critical to the sustainability of social media platforms and their ad economies.
 Minority groups, in particular, can face more privacy risks than others in a social network.
-For example, a few adversarial nodes can coordinate their actions to move a web crawler towards or away from certain parts of an online social network to compromise their privacy through a re-identification attack (Backstrom et al.
+For example, a few adversarial nodes can coordinate their actions to move a web crawler toward or away from certain parts of an online social network to compromise their privacy through a re-identification attack (Backstrom et al.
 2007).
-More broadly, deployment of learning systems to improve user experience requires accessing and aggregating historical usage information (e.g., for emojis and web domains) to optimize resource overhead for computation and communication, while respecting the personal and sensitive nature of the collected data (Apple Differential Privacy Team 2017, Erlingsson 2014b).
-5 Simulation Results
+More broadly, the deployment of learning systems to improve user experience requires accessing and aggregating historical usage information (e.g., emojis and web domains) to optimize resource overhead for computation and communication, while respecting the personal and sensitive nature of the collected data (Apple Differential Privacy Team 2017, Erlingsson 2014b).
+5 Simulations with Synthetic and Empirically-Grounded Network Data Sets
+We start with a sanity check using synthetic random graph models in Section 5.1 and then test our proposed framework for data collection and intervention design using highly sensitive data on sexual activity in an HIV prevention context in Section 5.2.
+5.1 Simulations on Synthetic Random Graph Models
+To test the performance of our algorithms in a sanitized environment, we generate an input using the Erdős-Rényi random graph model on $200$ nodes with edge probability $0.15$ and run the randomized response algorithm and the exponential mechanism with different privacy budgets ($\epsilon=0.1,0.5$, or $1$) and different numbers of influence samples ($m$ ranging from zero to $500$).
+The independent cascade probability is set to $0.03$.
+Figure 1 shows that the increase rate with $m$ slows with decreasing privacy budget for both the exponential mechanism and the randomized response algorithm.
 Figure 1: We show the performance (expected influence $I_{\mathcal{G}}(\cdot)$) of the outputs of our exponential mechanism and randomized response algorithms, as well as the non-private, influence sample seeding algorithm and the complete-information, deterministic greedy baseline over a synthetic network generated using Erdős-Rényi model with parameters $n=200$ and $p=0.15$.
@@
 Finally, to evaluate the performance of our mechanisms, we use another $1000$ influence samples to estimate the expected size of the spread.
-Simulations on a synthetic input graph.
-To test the performance of our algorithms in a sanitized setup, we generate an input using the Erdős-Rényi random graph model on $200$ nodes with edge probability $0.15$ and run our randomized response and exponential mechanism algorithms with different privacy budgets ($\epsilon=0.1,0.5$, or $1$) and different numbers of influence samples ($m$ ranging from zero to $500$).
-The independent cascade probability is set to $0.03$.
-Figure 1 shows that the rate of increase with $m$ slows down with the decreasing privacy budget for both exponential mechanism and randomized response algorithm alike.
-Simulations on a real-world dataset.
-We use a real dataset representing email interactions between individuals in a large European research institution taken from Stanford Large Network Dataset Collection (Stanford Network Analysis Project 2007, Yin et al.
-2017, Leskovec et al.
-2007).
-The nodes represent the individuals and the edges if there is communication between two individuals.
-This network represents the interaction between $1005$ nodes (individuals) through $25571$ edges.
-The interactions are mutual, and the graph is undirected.
-In the simulation setup of Fig.
-2, we consider $k=4$ and $k=8$ seeds, the number of influence samples ranges from zero to $1500$, $m\in\{0,100,250,500,750,1000,1250,1500\}$, and the independent cascade probabilities are set to $p=0.0155$ on every edges in ${\mathcal{G}}$ and zero otherwise.
-The $m=0$ case corresponds to no input information where the choice of $k=4,8$ seeds will be uniformly at random.
+5.2 Simulations on Empirically-Grounded Sexual Activity Networks
+We constructed a dynamic network of men who have sex with men (MSM) using separable temporal exponential random graph models (STERGM) based on data from cross-sectional egocentrically-sampled surveys in the ARTnet study (Anti-Retroviral Therapy Network), conducted in the United States from 2017 to 2019, with 4,909 participants, documenting 16,198 sexual partnerships (Weiss et al.
+2020).
+ARTnet collected data from individuals within the MSM target population, asking them to provide information on the quantity, characteristics, and history of their sexual relationships with other men.
+In this study, we focus on the San Francisco area due to the availability of data and its representative characteristics within the MSM population (San Francisco has a large and diverse MSM community with a long history of HIV-related public health interventions).
+Based on this data set, we estimate the STERGM parameters in Appendix 8.
+Using estimated STERGMs, we simulate sexual behavior among 1,000 individuals over 120 periods of one week each.
+Each node has fixed attributes, such as gender and age.
+We grouped every 12 periods as a quarter since our data is simulated on a weekly basis.
+For each quarter, we randomly selected 300 nodes as initial nodes.
+From these initial nodes, we identified all individuals connected directly or indirectly through sexual relationships.
+If node A had a direct sexual relationship with node B and node B subsequently had a relationship with node C, then nodes A and C are considered indirectly connected.
+This process resulted in 300 influence samples per quarter and a total of 3,000 influence samples.
+By selecting different initial nodes, we generated another collection of influence samples that are used to evaluate the seed sets that are selected by our algorithm.
+Unlike our previous approach, this time we no longer assume the existence of a latent network to construct influence samples through the independent cascade model.
+Instead, we focus directly on influence samples that better reflect the realities of contact tracing.
+In reality, it is typically easier to track sexual behavior than to obtain a complete network of new sexual relationships.
+Using influence samples derived from sexual behavior, our algorithm can identify high-risk nodes for targeted intervention, such as Pre-Exposure Prophylaxis (PrEP).
+PrEP is a strategy that uses antiretroviral drugs to reduce the risk of HIV infection for people who are not HIV positive but are at risk through sexual activity or injection drug use.
+Such interventions can be made cost-effective by focusing on individuals at high risk of HIV exposure Kessler et al.
+(2014), Juusola et al.
+(2012), Ross et al.
+(2016).
+The ARTnet data from 2017 to 2019 coincide with the PrEP uptake period, during which PrEP became increasingly accessible and widely adopted in the MSM community (in 2019, there were approximately 626,000 PrEP users in 77 countries, marking a 69% increase from 2018 (Schaefer et al.
+2021)).
+However, the data required to construct sexual activity or need-sharing networks is highly sensitive, making it crucial to implement proper privacy protections for individuals who share their data for public health interventions.
+We ran our privacy-preserving algorithm on influence samples constructed from ARTnet study data.
+In this case, the chosen seeds are high-risk nodes whose protection provides the most overall benefit in terms of preventing large disease cascades.
+In the simulation setup of Figure 2, we consider $k=10$ and $k=20$ seeds, the number of influence samples ranges from zero to $3000$, $m\in\{0,500,1000,1500,2000,2500,3000\}$.
+The case $m=0$ corresponds to the absence of input information where the choice of 10 or 20 seeds will be uniformly random.
 With no input information at $m=0$, there are no privacy risks to consider.
-Figure 2: We test the performance (expected influence $I_{\mathcal{G}}(\cdot)$) of our privacy-preserving seeding algorithms (exponential mechanism and randomized response) on a real-world dataset, at different privacy budgets ($\epsilon$), as number of influence samples ($m$) increases.
-The figure also shows the non-private, influence sample seeding and the complete-information, deterministic, greedy baselines for comparison.
-Each point on the plot is average of $50\times 10=500$ trials where we generate $50$ collections of $m$ influence samples, and run our mechanisms on each collection $10$ times.
-The $50$ realizations capture the randomness in the diffusion process (influence sampling) and for each set of influence samples, we perform $10$ iterations to average out the randomness of the algorithms (randomized response and exponential mechanism).
-The error bars show the $95\%$ confidence intervals of $500$ trials around the means.
-Finally, to evaluate the performance of our mechanisms, we use another $2000$ influence samples to estimate the expected size of the spread.
+Figure 2: We evaluate the performance of our privacy-preserving seeding algorithms (exponential mechanism and randomized response) on the MSM dataset by measuring their expected influence ($I_{\mathcal{G}}(\cdot)$) under different privacy budgets ($\epsilon$) as the number of influence samples ($m$) increases.
+For comparison, the plot also includes non-private baselines from seeding using influence samples and the deterministic greedy with complete network information.
+Each point on the plot represents the average of 1,000 trials, derived from 50 sets of $m$ influence samples, with each set subjected to 20 iterations of our randomized mechanisms.The error bars display 95% confidence intervals based on these 1,000 trials, but the intervals are too small to be visible.
+To assess the performance of our mechanisms in each case, we use a separate set of 3000 influence samples to estimate the expected spread size from the seed points that algorithms select.
 In both Figures 1 and 2, we see that the exponential mechanism outperforms the randomized response mechanism as the latter provides stronger (local) privacy protections.
-We also observe that low values of privacy budget $\epsilon$ lead to earlier saturation in lower quality expected spread sizes.
-In those cases, our maximum number of influence samples is not enough to counteract the random noise introduced by the privacy mechanisms, so the value of network information diminishes with the decreasing privacy budge — cf., Eckles et al.
+We also observe that low values of privacy budget $\epsilon$ lead to earlier saturation in lower expected spread sizes.
+In those cases, our maximum number of influence samples is not enough to counteract the random noise introduced by privacy mechanisms, so the value of network information diminishes with the decreasing privacy budge – cf.
+Eckles et al.
 (2022, Section 4).
@@
 In fact, the performance of both mechanisms approaches the original (“non-private”) influence sample greedy seeding algorithm (Algorithm 1), as $\epsilon\to\infty$.
-For instance, in the randomized response mechanism (Algorithm 4), as the privacy budget increases, the flip probability of the entries in the influence samples $x\in\{0,1\}^{n\times m}$ (Algorithm 4, line 3) decreases, and matrix $C_{\rho,l}$ in Equation 11 is close to the identity matrix, so the performance is similar to the original influence sample greedy seeding algorithm (Algorithm 1).
-For the exponential mechanism of Section 3.1, at each step the mechanism approximately chooses an optimal seed that has the largest expected marginal contribution to the spread, and a large privacy budget ($\epsilon\to\infty$) sharpens the probability of choosing an optimal seed against others.
-Figures 1 and 2 also show the deterministic greedy baseline with complete network information to be compared with the exponential mechanism and randomized response algorithms at different $\epsilon$ values with increasing $m$.
+For example, in the randomized response mechanism (Algorithm 4), as the privacy budget increases, the flip probability of the entries in the influence samples $x\in\{0,1\}^{n\times m}$ (Algorithm 4, line 3) decreases and the matrix $C_{\rho,l}$ in Equation 3 approaches the identity matrix, so the performance is similar to greedy non-private seeding on influence samples in Algorithm 1.
+For the exponential mechanism of Section 3.1, at each step the mechanism chooses the seeds probabilistically with probabilities proportional to the exponential of their marginal contribution to the spread size computed on the influence samples.
+In this case, increasing the privacy budget ($\epsilon\to\infty$) increases the probability of choosing the seed with the largest marginal contribution.
+Figures 1 and 2 also show the deterministic baseline for greedy with complete network information to be compared with the exponential mechanism and randomized response algorithms at different $\epsilon$ values with increasing $m$.
 As we gain better information about the network by increasing the number of influence samples $(m\to\infty)$ and the privacy budget $(\epsilon\to\infty)$, the performance of both algorithms will converge to the deterministic greedy algorithm with complete network information.
 6 Discussions, managerial insights and future directions
-Comparison between node-, edge-, and influence-sample DP.
-In this paper, we propose a new notion of differential privacy using influence samples adjacency.
-This notion may be extended to various problems which have underlying network structure.
-Examples include viral content and rumor cascades on social media, polls in elections from voters who influence each other’s opinions, and medical history of past disease contagions on epidemiological networks.
-Note that the input of an influence sample DP algorithm is different from the inputs of node- or edge-DP algorithms – capturing different notions of privacy loss that are not formally comparable; specifically, in influence sample DP an individual may decide whether or not to report their particular value for an influence sample but they cannot change the data generation process.
-Instead of entry-level adjacency, one may consider individual-level adjacency that two inputs are adjacent if they differ at one row (an individual can decide whether or not to report their entire contagion history).
-Under this new notion, our analysis for exponential mechanism in Section 3.1 still hold, but we need to strengthen privacy constraints from $\epsilon$ to $\epsilon/m$ where $m$ is the number of samples, because an individual’s change on the influence sample increases from $1$ to $m$.
-However, as the privacy constraints depend on the number of samples, a large number of samples can no longer ensure that our mechanisms’ performance converges to the original greedy algorithm with full knowledge of the graph.
+Comparison between node, edge, and influence-sample DP.
+In this paper, we propose a new notion of differential privacy based on the adjacency of influence samples.
+This concept can be extended to a wide range of problems that involve an underlying network structure.
+Examples include viral content and rumor cascades on social networks, polls in elections where voters influence each other’s opinions, and the medical history of past disease contagions in epidemiological networks.
+It is important to note that the input for an influence sample DP algorithm differs from that of node- or edge-DP algorithms, as they capture different notions of privacy loss that are not formally comparable.
+Specifically, in the influence sample DP, an individual can decide whether or not to report their particular value for an influence sample, but they cannot alter the data generation process.
+Sometimes, the underlying network may only exist as an abstract construct resulting from other physically defined relationships, such as the sexual activity network discussed in Section 5.
+In public health applications such as HIV prevention, where it is difficult to construct a complete sexual activity network due to privacy concerns, contact-tracing data from network cascades may be the only available input that can receive meaningful privacy protection.
+Our results address entry-level privacy by defining adjacency between cascade entries in the input data set.
+This can be generalized to individual-level privacy by considering that two inputs are adjacent if they differ in one column, allowing an individual to decide whether or not to report their entire contagion history.
+By focusing on influence samples rather than the entire network, we ensure that privacy concerns in sensitive contexts, such as HIV prevention, are addressed while still allowing effective data-driven interventions.
+Under this new notion, our analysis for the exponential mechanism in Section 3.1 still holds, but we need to reduce the privacy budget from $\epsilon$ to $\epsilon/m$, to account for the fact that each individual can modify the entire set of $m$ influence samples.
+However, as the privacy constraints depend on the number of samples, a large number of samples can no longer ensure that our mechanism’s performance converges to the original greedy algorithm with full knowledge of the graph.
 This is inevitable as the new notion prevents our mechanisms from finding high-quality seeds regardless of the number of samples.
-Indeed, it is impossible to get any meaningful utility for ISDP with individual-level adjacency for similar reasons as edge-DP or node-DP (LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg).
+In fact, it is impossible to get any meaningful utility from ISDP with individual-level adjacency for similar reasons as edge or node DP (LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg).
 Adjacency relation and differential privacy.
-In this paper, we see a separation of utility guarantee between node/edge-DP algorithms and influence sample DP algorithms (LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg vs LABEL:thm:exp:acc and LABEL:thm:rr:acc).
-Such separations may be attributed to different notions of adjacency (node, edge, or influence sample adjacency).
-This observation suggests that differentially private algorithms can have significantly different utility and privacy guarantees by changing the notion of adjacency.
-However, we are not advocating influence sample DP as superior to node- or edge-DP.
-On the contrary, one goal of this paper is to bring attention to the role that different notions of adjacency play in discussions of privacy.
-For instance, because people are connected by social networks, their private data of any forms (census, medical records, etc.) are inevitably affected by the network structure.
-Those supporting node- and edge-DP may consider that the proper notion of adjacency for two datasets (on census, medical records, etc.) should also depend on the network rather than the traditional Hamming distance.
+In this paper, we see a separation of utility guarantee between node/edge-DP algorithms and influence-sample DP algorithms (LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg versus LABEL:thm:exp:acc and LABEL:thm:rr:acc).
+Such separations may be attributed to different notions of adjacency (node, edge, or influence-sample adjacency).
+This observation suggests that DP algorithms can have significantly different utility and privacy guarantees by changing the notion of adjacency.
+Here, we are not advocating the influence sample DP as superior to node or edge DP.
+In contrast, we would like to bring attention to the role that different notions of adjacency play in discussions of privacy.
+For instance, because people are connected by social networks, their private data of any form (census, medical records, etc.) are inevitably affected by the network structure.
+Those supporting node and edge DP may consider that the proper notion of adjacency for two datasets should also depend on the network structure rather than the traditional Hamming distance between the entries.
 Further discussion can be found in Rezaei and Gao (2019).
@@
 To improve running times, one may consider improving the iterative sampler for the exponential mechanism in Algorithm 3, or designing a better post-processing step for Line 3 in Algorithm 5.
-On the other hand, our ISDP randomized response algorithm $\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}})$ is automatically robust against independent randomness noise on influence samples.
-However, it is unclear how well our ISDP exponential mechanism $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ performs in the presence of noise.
+On the other hand, our ISDP randomized response algorithm $\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}})$ is automatically robust against independent random noise in influence samples.
+In contrast, it is unclear how well our ISDP exponential mechanism $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ performs in the presence of noise.
 Conclusions.
-The correlated nature of social network information has been noted as a vulnerability in DP formulations which provide user-level guarantees with respect to the addition or removal of a single user’s data but ignore information leakage across different users’ records.
-By focusing on the utility of influence samples for influence maximization, our approach guarantees user’s participation in any given cascade is kept private either locally or centrally and preserves the statistical validity of the cascades for influence estimation and maximization.
-Subsequently, we provide a rigorous operationalization of action-level privacy by providing users with plausible deniability as to whether or not they have taken part in any particular network cascade (Kifer et al.
+The correlated nature of social network information is a vulnerability in DP formulations that provide user-level guarantees with respect to the addition or removal of a single user’s data, but ignore information leakage across different user records.
+By focusing on the utility of influence samples for influence maximization, our approach guarantees user participation in any given cascade to be kept private either locally or centrally, while preserving the statistical validity of the cascades for influence estimation and maximization.
+Subsequently, we provide a rigorous operationalization of action-level privacy by providing users with plausible deniability as to whether or not they have participated in any particular network cascade (Kifer et al.
 2020).
-This relaxation of user-level DP is ideally suited for the purposes of data collection to inform network intervention design.
-We expect that formulation and investigation of influence sample DP, with specific attention to the utility of the collected data for network intervention design, should lead to novel generalizations of privacy that are both consequentialist and robust to variations in sampling distributions and randomization noise (Kifer and Machanavajjhala 2014, Rastogi et al.
+This relaxation of user-level DP is ideally suited for contact tracing purposes to inform public health interventions, such as with the sexual activity network presented in our simulation results.
+We expect that formulation and investigation of the influence sample DP, with specific attention to the utility of the collected data for the design of network intervention, should lead to novel generalizations of privacy that are consequentialist and robust to variations in sampling distributions and randomization noise (Kifer and Machanavajjhala 2014, Rastogi et al.
 2009, Liu et al.
@@
 2022).
-[…]
-Rahimian and Yu contributed equally and are listed alphabetically.
-Rahimian is partially supported by the National Science Foundation under Grant No.
-2318844.
-Rahimian also acknowledges support from a Pitt Momentum Funds Grant on Socially Responsible Data Collection and Network Intervention Designs.
-This research was supported in part by the University of Pittsburgh Center for Research Computing, RRID:SCR_022735, through the resources provided.
-Specifically, this work used the H2P cluster, which is supported by NSF award number OAC-2117681.
-Code for reported simulations can be accessed from https://github.com/aminrahimian/dp-inf-max
+Data Availability
+The data used to generate the MSM dynamic sexual behavior dataset originates from the ARTnet study, accessible at https://github.com/EpiModel/ARTnet.
+Due to confidentiality concerns, this dataset is not publicly accessible.
+Access to the data requires a Memorandum of Understanding (MOU), which specifies the personnel involved in the analysis and the purposes of the data analysis.
+The dataset cannot be shared without the consent of the ARTnet Study Principal Investigator (PI), Samuel Jenness at Emory University, as outlined in the MOU.
+To request access, please contact the PI via email at samuel.m.jenness@emory.edu.
+Upon review and approval of the MOU, access to the ARTnet dataset will be provided through Github.